### v1

# ComGPT: Detecting Local Community Structure with Large Language Models

Abstract
Large language models (LLMs), like GPT, have demonstrated the ability to understand graph structures and have achieved excellent performance in various graph reasoning tasks such as node classification. So far, how to leverage LLMs to better detect local communities remains underexplored. Local community detection algorithms based on seed expansion often face a seed-dependent problem, community diffusion, and free rider effect. Using LLMs to solve existing local community work problems faces the following challenges: existing graph encoding methods fail to provide LLMs with sufficient community-related graph information; LLMs lack domain knowledge in mining communities. To address these issues, we improve graph encoding by supplementing community knowledge to enhance the ability of graph encoding to express graph information. Additionally, we design the NSG (Node Selection Guide) prompt to enhance LLMs’ understanding of community characteristics, aiming to alleviate the seed-dependent problem, community diffusion, and free rider effect. Based on the graph encoding and NSG prompt, we present a GPT-guided local community detection, called ComGPT. ComGPT iteratively selects potential nodes from the detected community’s neighbors and subsequently employs GPT to choose the node that optimally integrates into the detected community from these selected potential nodes. Experimental results demonstrate that ComGPT outperforms the comparison algorithms, thereby confirming the effectiveness of the designed graph encoding and prompt.
Introduction
Identifying community structures from networks has wide applications, such as identifying protein complexes (Yin et al. 2017). Local community detection has gained attention for its ability to quickly identify the community containing a specific node (Ni et al. 2024a, b). Various methods are designed for detecting local community, including seed expansion algorithms (Luo, Wang, and Promislow 2008), heat diffusion techniques (Kloster and Gleich 2014), $k$-core decomposition (Cui et al. 2014), $k$-truss identification (Faysal et al. 2023), and non-negative matrix factorization algorithms (Wang et al. 2011). Among these methods, the seed expansion algorithm has garnered widespread acclaim for its versatility and ease of implementation. The seed expansion algorithms typically enlarge the community by selecting a node with the highest quality function value at each step. However, it often faces the seed-dependent problem (Kloumann and Kleinberg 2014), community diffusion, and free rider effect (Sozio and Gionis 2010; Lang and Andersen 2007; Tsourakakis et al. 2013). To address these problems, researchers developed different strategies such as identifying core node sets and designing membership functions with a bias towards nodes with large degrees (Luo et al. 2018).
In this study, we explore an alternative approach inspired by the the success of large language models (LLMs) in many fields (Gu et al. 2024; Zong and Krishnamachari 2024; Yang et al. 2022) such as theory-of-mind reasoning (Adhikari et al. 2020; Ammanabrolu and Riedl 2021) and graphical reasoning (Wang et al. 2023). As far as we know, LLMs have not been applied to local community detection. Therefore, we explore the application of large models for local community detection and propose an algorithm, called ComGPT (GPT-guided Local Community Detection). This algorithm leverages GPT to guide the detection process, thereby bridging the current gap in LLMs applicability to this field. ComGPT integrates the seed expansion algorithm with GPT, leveraging their inherent compatibility: GPT through step-by-step guidance, while the seed expansion method incrementally adds nodes to the community, making their integration highly synergistic.
However, the application of LLMs to local community detection faces the following challenges: 1) Lack of graph encoding methods for providing LLMs with sufficient community-related graph information. To address this, we enhance graph encoding by incorporating two key aspects: graph topology and community information. 2) Lacks domain knowledge for LLMs in dealing with local community detection. It limits the capabilities of LLMs. To address this, we have designed prompt to assist LLMs in understanding community definitions and how to select nodes. 3) Lacks efficient interaction between GPT and community detection methods. The mode for integrating community detection algorithms with GPT, as well as strategies for reducing the number of input tokens during their interaction, remain unexplored. Based on the step-by-step characteristics of GPT and seed expansion algorithm, we design a step-by-step interactive mode. To minimize costs by reducing the number of tokens, the input to GPT is not the entire graph but a partial graph that includes the detected community and some of its neighbors. Our contribution is summarized as follows:
  • We propose a GPT-guided local community detection algorithm, called ComGPT. ComGPT repeatedly selects potential nodes from the neighbors of the detected community and then uses GPT to identify the node that optimally integrates into the detected community from these selected potential nodes. To the best of our knowledge, this is the first work to apply LLMs to the domain of local community detection.
  • For GPT to understand and detect communities, we design graph encoding method, NSG prompt, and interaction mode. The graph encoding method incorporates both graph topology and community information, which enhances GPT to understand the current state of the detected community. NSG prompt that integrates community features facilitates GPT’s comprehension of domain knowledge. Additionally, an interactive mode is established between GPT and the local community detection method.
  • The results on real datasets indicate that ComGPT outperforms comparison methods. Ablation study confirm that GPT plays a crucial role in the ComGPT ’s performance. Furthermore, prompt comparisons experiments reveal that the designed prompt effectively aids GPT in understanding the graph structure and community situation.
Related work
Local community detection
Local community detection focuses on discovering the community to which the seed node belongs (Ni et al. 2024c, 2023). A variety of methods like seed expansion algorithms are developed for detecting local communities (Luo, Wang, and Promislow 2008; Cui et al. 2014), which expand communities by selecting nodes with the highest quality function values. For example, the R method (Clauset 2005) and M method (Luo, Wang, and Promislow 2008) choose nodes based on the highest local modularity R and M, respectively. The seed expansion algorithms often faces the seed-dependent problem (Kloumann and Kleinberg 2014), community diffusion, and free rider effect (Sozio and Gionis 2010; Lang and Andersen 2007; Tsourakakis et al. 2013). To address the seed-dependent problem, researchers developed different strategies such as identifying core node sets (Guo et al. 2022), finding the nearest nodes with greater centrality (Luo et al. 2020), and designing membership functions with a bias towards nodes with large degrees (Luo et al. 2018). Our work mitigates the aforementioned problems by utilizing GPT, distinguishing it from existing studies.
LLMs
In recent years, researchers have explored LLMs for node classification (Chen et al. 2023) and graph reasoning (Zhang 2023). Wang et al. demonstrated LLMs’ graph reasoning abilities and designed the NLGraph dataset (Wang et al. 2023) to evaluate these capabilities. Furthermore, Guo et al. emphasized the design and improvement of graph encoding techniques, which critically affect GPT’s interpretation of input data and its subsequent output (Guo et al. 2023). Therefore, Adjacency and Incident (Fatemi, Halcrow, and Perozzi 2023) are designed to encode graph structured data into text for LLMs. Adjacency uses integer node encoding and parenthesis edge encoding. Incident uses integer node encoding and incident edge encoding. Existing graph coding methods primarily represent graph topology, which is often insufficient for specific tasks, such as community detection. Our work introduces a novel graph encoding method tailored for the community detection domain, which incorporates community structure considerations, distinguishing it from existing approaches.
Furthermore, various prompts are designed for specific tasks to make LLMs better adapt to these tasks (Sahoo et al. 2024), such as the general COT prompt (Wei et al. 2022) and the BAG prompt (Wang et al. 2023) for graph reasoning tasks. However, these prompts do not incorporate the domain knowledge required for local community detection and are unsuitable for detecting community structures within networks. This paper designs a prompt, named NSG, suitable for the field of local community detection, which is different from the above work.
Proposed algorithm
Shortcomings
Existing heuristic algorithms suffer from the seed-dependent problem (Kloumann and Kleinberg 2014), community diffusion, and free rider effect (Sozio and Gionis 2010) (Lang and Andersen 2007) (Tsourakakis et al. 2013), which are explained as following.
(a) Seed-dependent problem
(b) Community diffusion
(c) Free rider effect
Figure 1: Existing shortcomings. Nodes in the same community are marked with the same color.
Shortcoming 1: Seed-dependent problem (Kloumann and Kleinberg 2014).
It refers to the quality of the community detected being highly dependent on the initial seed nodes, as illustrated by nodes 5 and 1 in Figure 1(1(a)). The communities detected by the M method (Luo, Wang, and Promislow 2008) with nodes 5 and 1 as the starting seed nodes are [5, 11, 6, 7, 17] and [1, 12, 13, 4, 8, 2, 14, 3, 22, 18, 20], respectively. The former includes some nodes of the real community and misses many nodes, while the latter is more accurate.
Shortcoming 2: Community diffusion.
The community diffusion problem means that the detected community contains nodes from different communities. Take node 1 in Figure 1(1(b)) as an example to illustrate the community diffusion problem. The community detected by the M method with node 1 as the starting seed node is [1, 2, 3, 4, 7, 8, 9, 13, 14, 15, 16, 17, 18, 20], which includes nodes from two different communities.
Shortcoming 3: Free rider effect (Sozio and Gionis 2010) (Lang and Andersen 2007) (Tsourakakis et al. 2013).
The free rider effect refers to mistakenly adding unrelated communities to enhance the fitness of a detected community. We use the community derived from seed node 7, as shown in Figure 1(1(c)), to illustrate it. The target community consists of nodes 5, 6, 7, 8, and 9. However, to achieve a higher fitness value, the detected community not only contains nodes in the target community [5, 6, 7, 8, 9] but also erroneously includes all nodes in the unrelated community [1, 2, 3, 4, 10, 11].
ComGPT
Figure 2: The diagram of ComGPT.
We propose a local community detection algorithm based on GPT, called ComGPT. Figure 2 shows the main steps of ComGPT. It includes four main steps: 1) potential nodes identification; 2) graph encoding; 3) node selection by GPT; 4) node supplementation by GPT. Potential nodes identification aims to pinpoint potential nodes from the neighbors of the current community. Graph encoding involves transforming the graph structure into a textual format interpretable by GPT. Node selection is that GPT selects the node that optimally integrates into the community from potential nodes. Node supplement involves GPT selecting which nodes to add to the community when no potential nodes remain.
Algorithm 1 ComGPT
Input: Network $G$, Seed node $v_{s}$
Output: Community $C$
1:  $C\leftarrow{\{v_{s}\}}$, $cands\leftarrow\emptyset$
2:  while True do
3:     $PN\leftarrow$ potential nodes identification with $C$
4:     while $PN\neq\emptyset$ do
5:        $gtxt\leftarrow$ graph encoding with $C$ and $N$
6:        $v\leftarrow$ node selection by GPT from $PN$
7:        add $v$ to $C$
8:     end while
9:     add $C$ to $cands$
10:     $v\leftarrow$ node supplementation by GPT
11:     if $v$ is not null then
12:        add $v$ to $C$
13:     else
14:        break
15:     end if
16:  end while
17:  $C\leftarrow$ community with highest $M$ in $cands$
18:  return  $C$
Algorithm 1 shows the overall process of ComGPT. For convenience, we briefly introduce the notations used in this paper. Let $G$ denote the network, $v_{s}$ denote the seed node, $C$ denote the detected community, $v$ denote the node returned by the node selection or node appending steps, $gtxt$ denote the text of the graph structure obtained through graph encoding, $PN$ denote the set of potential nodes, $cands$ store the community each time $PN$ is empty. Initially, community $C$ consists of only one seed node $v_{s}$ and the $cands$ is empty (line 1). First, ComGPT selects potential nodes $PN$ from neighbors of current community $C$ for GPT (line 3). The local network, i.e., the subgraph consisting of the nodes of the current community $C$ and $PN$, is converted into text, i.e., $gtxt$ (line 5). Utilizing $gtxt$ and NSG (Node Select Guide) prompt, GPT selects the node $v$ from potential nodes $PN$ and adds it to the current community $C$ (line 6,7). The above steps are repeated until $PN$ is empty. After the potential nodes $PN$ is empty, the community $C$ is added to the candidate communities $cands$ (line 9). When $PN$ is empty, it does not necessarily imply that the neighboring nodes of the community are weakly connected to the community. Instead, it could be that the nodes have relatively high degrees, which marginally reduces the $M$ value. To address this scenario, ComGPT identifies nodes through node supplementation (line 10). If GPT identifies suitable node $v$ to add to the community, then ComGPT adds $v$ to $C$ and continues to expand current community $C$ (line 12). Otherwise, ComGPT stops community expansion (line 14). After the above process is completed, ComGPT selects the community with the highest $M$ value from $cands$ as the final community (line 17).
Potential nodes identification
To mitigate the high computational cost associated with processing nodes through GPT, ComGPT identifies potential nodes that may belong to the current community and provides these nodes to GPT for selection. Different fitness functions are designed in the seed expansion algorithm to select nodes, which can be used to choose potential nodes, such as modularity $M$ (Luo, Wang, and Promislow 2008), $R$ (Clauset 2005), and $Q_{lcd}$ (Shang et al. 2023). ComGPT uses local modularity $M_{C}$ to select potential nodes, which is defined as the ratio of the number of internal edges in community $C$ to the number of external edges of community $C$. Nodes that are satisfied $\Delta M=M_{C\cup\{v\}}-M_{C}>0$ are regarded as potential nodes. During the initial expansion of the community, since nodes with high degrees often struggle to achieve $\Delta M>0$, these nodes are not considered potential nodes $PN$. To alleviate this preference, these nodes are added to potential nodes based on their closeness value to the current community $C$. The closeness value $f_{v}$ between node $v$ and the community $C$ is calculated as
$f_{v}=\frac{|C\cap N(v)|}{|C|},$ (1)
$|C|$ represents the number of nodes in the community and $N(v)$ represents the neighbors of node $v$. The value of $f_{v}$ ranges from 0 to 1. Higher values indicate a greater proportion of nodes connected to those within the community.
The initial potential nodes consist of nodes with $\Delta M>0$ and nodes with high $f_{v}$ values. To further reduce the number of nodes processed by GPT, $k$ nodes are selected from them to form a potential node’s set.
Graph encoding
Figure 3: Graph encoding.
The purpose of graph encoding is to convert graph structure data and information about the community $C$ into text that is easily understandable by GPT. The graph data includes two aspects. 1) Graph topological structure. We adopt the Incident method (Fatemi, Halcrow, and Perozzi 2023) to let GPT understand the graph’s topological structure. The Incident method represents the graph topology by introducing the first-order neighbors of each node. To reduce computational overhead, the graph structure inputted to GPT is a local network, i.e., the subgraph consisting of the nodes of the current community $C$ and $PN$. For clarity, the middle part of Figure 3 illustrates the specific graph encoding: “G describes a graph among nodes:1, 2, 3, 5, 6, 7. Node 1 is connected to nodes 2,3”. To reduce the number of tokens, some conjunctions that do not affect the understanding of GPT are removed. 2) Supplementary knowledge. We describe node information from a community perspective to enhance GPT’s understanding of community structure. This involves outlining which nodes are within the community, those that are outside, and detailing their specific external connections. For clarity, the lower part of Figure 3 illustrates graph encoding for community $C$ = [1, 2, 3] and potential nodes $PN$ = [5, 6, 7]: “Nodes in the current community:[1,2,3]. The outside nodes contain:[5,6,7]. Node 5 is connected to nodes in community:[1,3]”.
Node selection by GPT
Node selection refers to choosing the most suitable node for community expansion, based on the prompt and graph data. Below, we introduce node selection through prompts that teach GPT to accurately select nodes.
1. The more a potential node is connected to other potential nodes, the higher the likelihood of its selection: This prompt could alleviate the seed-dependent problem and the free rider effect.
The reason for the seed-dependent problem is low-degree node preference. Take seed node 5 in Figure 1(1(a)) and M method (Luo, Wang, and Promislow 2008) as an example to illustrate it. Nodes 1, 7, and 11 are each connected to the node in the community [5] by an edge, and node 11, having the smallest degree, is selected so that the community has the fewest external edges. Thus, the M method selects node 11 to join the community. Nodes that connect to many other potential nodes are usually high-degree nodes, which help to avoid bias towards low-degree nodes and thereby alleviate the seed-dependent problem. Taking Figure 1(1(a)) as an example, the community includes node 5, with nodes 1, 7, and 11 regarded as potential nodes. According to prompt 1, node 1 will be selected over the mistakenly considered node 11.
The reason for the free rider effect is the neglect of the connectivity of neighbor nodes of the detected community. Nodes that connect to many other neighbor nodes are consideration of their association with other potential nodes, which alleviates the free rider effect. Taking Figure 1(1(c)) as an example, the community contains node 7, with nodes 4, 6, and 8 regarded as potential nodes. According to prompt 1, nodes 6 and 8 have a higher probability of being selected. Node 4 is a node that may cause the free rider effect, and will not be selected.
2. Prioritize selecting nodes that are connected to multiple nodes within the community: This prompt could alleviate the community diffusion problem. The reason for the community diffusion is the insufficient consideration of the connections between nodes to be added to the community and the nodes within the community. Therefore, the prompt considers the connection between potential nodes and the community, thereby alleviating the community diffusion problem. Taking Figure 1(1(b)) as an example, the community includes nodes 1 and 3, with nodes 2, 5, 7, 13, 15, and 18 regarded as potential nodes, where node 13 or 18 may lead to community diffusion. According to prompt 2, nodes 2 and 5 have a higher probability of being selected, rather than mistakenly selecting nodes 13 or 18.
The first prompt is to consider the closeness with neighboring nodes outside the community, and the second prompt is to consider the closeness with nodes within the community. This is consistent with the characteristics of the community being close inside and sparse outside, thus helping GPT understand the characteristics of the nodes to be selected.
Node supplementation by GPT
When $PN$ is empty, ComGPT uses GPT to supplement nodes. Node supplementation involves evaluating whether to continue the expansion process or terminate it. If continuation is warranted, it also determines which additional nodes should be selected. In terms of graph encoding, the graph encoding method mentioned above, including graph topological structure and supplementary knowledge, is used. The difference is that to provide GPT with sufficient information to make a decision, the input graph text corresponds to graph topology that includes the current community and its first and second-order neighbors. The nodes to be selected by GPT are the three nodes with the largest M values among the first-order neighbor nodes of the community. In terms of prompt, it is based on the definition of the community and whether these nodes can maintain the cohesion of the community. If GPT determines that any of these three nodes warrant inclusion in the community, it returns the node’s number; otherwise, it returns NULL.
Experiments
To evaluate the performance of our algorithm, we conducted some experiments on real datasets. We first outline the experimental settings, which include datasets, comparison algorithms, and evaluation metrics. We then present the experimental results and provide corresponding analyses.
Experimental settings
dataset M R DMF_M DMF_R LCDNN LCDMD ComGPT
Dolphins 0.4685 0.4505 0.7255 0.7242 0.7558 0.7235 0.7885
Polbooks 0.5257 0.5030 0.7295 0.7286 0.7463 0.7301 0.7586
Football 0.7146 0.7107 0.8881 0.8893 0.8948 0.8134 0.8406
Amazon 0.7706 0.7554 0.8353 0.8361 0.8372 - 0.8385
DBLP 0.4185 0.4185 0.4032 0.4009 0.3152 - 0.4247
Table 1: Comparisons with other methods in terms of fscore.
Datasets
Five real-world datasets, including Dolphins (Lusseau 2003), Football (Girvan and Newman 2002), Polbooks (Krebs 2004), Amazon (Yang and Leskovec 2012), and DBLP (Yang and Leskovec 2012) are utilized to test the performance of algorithms.
The Dolphins network is based on observations of associations between pairs of dolphins, where edges represent associations between dolphins. It contains 62 nodes and 159 edges. The Football network contains 115 nodes and 613 edges. Nodes within this network represent college football teams, communities correspond to different leagues, and edges correspond to games played between them. The Polbooks network is a co-purchasing network for books. Each node in the dataset represents a book about US politics. An edge between two books indicates that they are often purchased together by customers. It contains 105 nodes and 441 edges.The Amazon network represents product co-purchasing relationships, consisting of 334,863 nodes and 925,872 edges. In this network, nodes representing products and edges indicate which products are frequently co-purchased. The DBLP network is a co-authorship network comprising 317,080 nodes and 1,049,866 edges. Here, nodes represent authors, and edges indicate that two authors have co-authored a paper.
Comparison algorithms
To evaluate the performance of ComGPT, we compare it with five local community detection algorithms: M method (Luo, Wang, and Promislow 2008), R method (Clauset 2005), DMF_M (Luo et al. 2018), DMF_R (Luo et al. 2018), LCDNN (Luo et al. 2020), and LCDMD (Guo et al. 2022). ComGPT uses gpt-3.5-turbo-0125 and sets $k_{f}$ to 0.5. Due to monetary costs of GPT-3.5-turbo-0125, we execute ComGPT twice for each starting seed node to handle GPT hallucinations.
Evaluation metrics
The evaluation metrics precision, recall, fscore (Powers 2020) are used to evaluate the detected communities. For an overlapped seed node, all the communities it belongs to are integrated as a real local community. fscore is the harmonic mean of precision and recall. The values of precision, recall, and fscore are between 0 and 1. The larger value implies a better algorithm performance.
Experimental results
Results on networks
For the Dolphins, Polbooks, and Football datasets, each node is selected as a seed node separately. For the Amazon and DBLP datasets, We randomly select 200 nodes from the top 5,000 communities. The average fscore of all seed nodes is calculated. The detailed experimental results are shown in Table 1. Due to the unavailability of the LCDMD (Guo et al. 2022) source code, results for the Dolphins, Polbooks, and Football datasets are cited from (Guo et al. 2022), in which results for the Amazon and DBLP datasets are not available.
Table 1 shows that ComGPT outperforms the M method, R method, DMF_M, DMF_R, LCDNN, and LCDMD on most datasets. ComGPT is superior to other methods because GPT has excellent comprehension to find the suitable node based on the NSG prompt. On the Football dataset, ComGPT performs worse than DMF_M, DMF_R, and LCDNN. It is because Football contains several tightly interconnected communities, which leads GPT to erroneously identify them as a single community. Except for the DBLP dataset, DMF_M, DMF_R, LCDNN, and LCDMD perform better than the M method and R method. The reason why DMF_M, DMF_R, LCDNN, and LCDMD outperform the M method and R method on most datasets is that the latter two methods maximize local modularity (M and R), making them sensitive to the choice of seed nodes, known as seed-dependent problem. In contrast, DMF_M, DMF_R, LCDNN, and LCDMD implement various strategies to alleviate this problem. Specifically, DMF_M, and DMF_R, utilize dynamic membership functions to select nodes, which tend to favor nodes with large degrees. LCDNN expands the community based on the fuzzy relationships between nodes and NGC nodes. Meanwhile, LCDMD introduces a new modularity density to effectively mine the core parts of a community.
Ablation study
Method GPT SK NSG precision recall fscore
1 NO NO NO 0.6824 0.7246 0.6978
2 YES NO YES 0.7571 0.8277 0.7832
3 YES YES NO 0.5911 0.7026 0.6221
ComGPT YES YES YES 0.8440 0.8520 0.8445
Table 2: Results of ablation experiment on the football dataset.
To explore the necessity of GPT, SK (Supplementary knowledge in graph coding), and NSG prompt in ComGPT, we conduct ablation experiments on the Football dataset with 50 randomly selected seed nodes. Three comparative models are constructed as follows: 1) Method 1 does not use GPT. The first node in $PN$ will be selected for community expansion until $PN$ is empty; 2) Method 2 only maintains GPT and NSG prompt, without SK; and 3) Method 3 only maintains GPT and SK, without NSG prompt. When calculating precision and recall, it is necessary to obtain the intersection between the detected community, which may include duplicate nodes, and the true community. In this calculation, duplicate nodes in the detected community are regarded as a single node. To prevent cyclic selection of duplicate nodes, community expansion is terminated once the number of duplicate nodes exceeds one-third of the number of nodes in the community.
Table 2 shows the experimental results of Methods 1, 2, 3, and ComGPT. Overall, ComGPT outperforms Methods 1, 2, and 3, indicating that GPT, SK, and NSG prompt play a crucial role in ComGPT, respectively. ComGPT is better than Method 1 in terms of precision, recall, and fscore, which demonstrates the effectiveness and importance of GPT. The reason ComGPT outperforms Method 2 is that its graph coding incorporates community knowledge. Compared with graph encoding methods that contain only the graph topology, supplementing community knowledge enables the GPT to better understand the community structure and thus make more accurate judgments. ComGPT also outperforms Method 3, suggesting that NSG prompt can significantly enhance its performance in local community detection.
Prompt comparisons
 Method Dolphins Polbooks Football
Zero-shot P 0.7195 0.5793 0.5911
R 0.3911 0.2330 0.7026
F 0.4690 0.3261 0.6221
Few-shot P 0.7136 0.5658 0.6592
R 0.3840 0.3358 0.7600
F 0.4556 0.4070 0.6894
COT P 0.7483 0.5924 0.6120
R 0.4397 0.2633 0.7205
F 0.5173 0.3572 0.6512
BAG P 0.7242 0.6005 0.6530
R 0.4704 0.2338 0.7589
F 0.5192 0.3315 0.6870
NSG P 0.9592 0.7788 0.8440
R 0.7356 0.7559 0.8520
F 0.7964 0.7506 0.8445
Table 3: Results of different prompts. P, R, and F are the abbreviations of recall, precision, and fscore, respectively.
To test the performance of the NSG prompt within ComGPT, we replaced the NSG prompt with four typical prompt approaches: Zero-shot (Brown et al. 2020), Few-shot (Brown et al. 2020), COT (Wei et al. 2022), and BAG (Wang et al. 2023) prompts. These alternate versions of the algorithm serve as baseline models for comparison. We conducted comparative experiments by randomly selecting 50 seed nodes in the Polbooks, Dolphins, and Football datasets. The detected communities may contain duplicate nodes, and the indicator values are calculated in the same way as in the ablation study.
Table 3 shows that the performance of NSG is better than that of Zero-shot, Few-shot, COT, and BAG prompts. Specifically, on Polbooks and Football datasets, the Few-shot outperforms the Zero-shot. Few-shot is worse than Zero-shot on the Dolphins dataset. This is because each dataset has a different structure. If the provided examples match the community structure characteristics of the dataset, the performance of GPT is improved. Conversely, if they do not match, GPT’s performance is diminished. The reason COT and BAG outperform Zero-shot on most datasets is that COT enhances GPT’s reasoning ability by guiding it to think step-by-step and BAG helps GPT better digest graph information. On Dolphins, Polbooks, and Football datasets, NSG outperforms other prompts. This is because the design of COT and BAG is oriented towards universal graph reasoning problems so that no relevant guidance or knowledge is utilized in dealing with local community detection. NSG is oriented to the field of local community detection and utilizes the guidance in local community detection.
Effect of the number of potential nodes on ComGPT
Figure 4: Effect of $k$ on ComGPT.
We explore the effect of the number of potential nodes, denoted by $k$, on ComGPT on the Football dataset, with 50 seed nodes selected randomly. The values of $k$ are 5, 7, 10, 13, and 15. Figure 4 shows the performance of ComGPT under different $k$. As the $k$ increases, the performance of ComGPT remains generally stable, though with some fluctuations. Fluctuations are relatively low across different values, except for $k$ = 13, which exhibits higher variability due to the randomness of GPT output. To confirm this, we conduct repeated experiments for $k$ = 13 and the results are nearly identical to those obtained with other values of $k$. The above analysis suggests that the performance of ComGPT is less affected by the parameter $k$.
Conclusion
We explored the potential of using GPT to assist in local community detection and developed a GPT-guided methodology, termed ComGPT. Additionally, a graph encoding method suitable for local community detection is explored, which enhances the existing graph encoding by incorporating community knowledge. Furthermore, the NSG prompt is designed to enhance LLMs’ understanding of community domain knowledge. Experimental results demonstrate that ComGPT outperforms the comparison algorithms. Moreover, the designed graph encoding and prompt help GPT understand the domain knowledge of graph topology and community discovery, respectively.

### v2

@@
 Abstract
-Large language models (LLMs), like GPT, have demonstrated the ability to understand graph structures and have achieved excellent performance in various graph reasoning tasks such as node classification.
-So far, how to leverage LLMs to better detect local communities remains underexplored.
-Local community detection algorithms based on seed expansion often face a seed-dependent problem, community diffusion, and free rider effect.
-Using LLMs to solve existing local community work problems faces the following challenges: existing graph encoding methods fail to provide LLMs with sufficient community-related graph information; LLMs lack domain knowledge in mining communities.
-To address these issues, we improve graph encoding by supplementing community knowledge to enhance the ability of graph encoding to express graph information.
-Additionally, we design the NSG (Node Selection Guide) prompt to enhance LLMs’ understanding of community characteristics, aiming to alleviate the seed-dependent problem, community diffusion, and free rider effect.
-Based on the graph encoding and NSG prompt, we present a GPT-guided local community detection, called ComGPT.
-ComGPT iteratively selects potential nodes from the detected community’s neighbors and subsequently employs GPT to choose the node that optimally integrates into the detected community from these selected potential nodes.
-Experimental results demonstrate that ComGPT outperforms the comparison algorithms, thereby confirming the effectiveness of the designed graph encoding and prompt.
+Large Language Models (LLMs), like GPT, have demonstrated the ability to understand graph structures and have achieved excellent performance in various graph reasoning tasks, such as node classification.
+Despite their strong abilities in graph reasoning tasks, they lack specific domain knowledge and have a weaker understanding of community-related graph information, which hinders their capabilities in the community detection task.
+Moreover, local community detection algorithms based on seed expansion, referred to as seed expansion algorithms, often face the seed-dependent problem, community diffusion, and free rider effect.
+To use LLMs to overcome the above shortcomings, we explore a GPT-guided seed expansion algorithm named ComGPT.
+ComGPT iteratively selects potential nodes by local modularity $M$ from the detected community’s neighbors, and subsequently employs LLMs to choose the node to join the detected community from these selected potential nodes.
+To address the above issues faced by LLMs, we improve graph encoding method, called Incident, by incorporating community knowledge to improve LLMs’s understanding of community-related graph information.
+Additionally, we design the NSG (Node Selection Guide) prompt to enhance LLMs’ understanding of community characteristics.
+Experimental results demonstrate that ComGPT outperforms the comparison methods, thereby confirming the effectiveness of the improved graph encoding method and prompts.
 Introduction
@@
 2017).
-Local community detection has gained attention for its ability to quickly identify the community containing a specific node (Ni et al.
+Local community detection has gained attention due to its ability to quickly identify the community containing a specific node (Ni et al.
 2024a, b).
-Various methods are designed for detecting local community, including seed expansion algorithms (Luo, Wang, and Promislow 2008), heat diffusion techniques (Kloster and Gleich 2014), $k$-core decomposition (Cui et al.
-2014), $k$-truss identification (Faysal et al.
-2023), and non-negative matrix factorization algorithms (Wang et al.
+Various methods are designed for detecting local community, including seed expansion algorithms (Luo, Wang, and Promislow 2008), heat diffusion techniques (Kloster and Gleich 2014), and non-negative matrix factorization algorithms (Wang et al.
 2011).
-Among these methods, the seed expansion algorithm has garnered widespread acclaim for its versatility and ease of implementation.
-The seed expansion algorithms typically enlarge the community by selecting a node with the highest quality function value at each step.
-However, it often faces the seed-dependent problem (Kloumann and Kleinberg 2014), community diffusion, and free rider effect (Sozio and Gionis 2010; Lang and Andersen 2007; Tsourakakis et al.
-2013).
-To address these problems, researchers developed different strategies such as identifying core node sets and designing membership functions with a bias towards nodes with large degrees (Luo et al.
+Among these methods, the seed expansion algorithms have garnered widespread acclaim for their versatility and ease of implementation.
+The seed expansion algorithms typically enlarge the community by selecting a node with the high scoring function value at each step (Moradi, Olovsson, and Tsigas 2014).
+However, seed expansion algorithms often face the seed-dependent problem (Kloumann and Kleinberg 2014), community diffusion, and free rider effect (Sozio and Gionis 2010; Lang and Andersen 2007; Tsourakakis et al.
+2013), shown in Figure 1.
+The seed-dependent problem refers to the quality of the community detected being highly dependent on the initial seed nodes, as illustrated by nodes 5 and 1 in Figure 1(a).
+The community diffusion problem means that the detected community contains nodes from different communities, as illustrated by node 1 in Figure 1(b).
+The free rider effect refers to mistakenly adding unrelated communities to enhance the scoring function value of a detected community, as illustrated by node 7 in Figure 1(c).
+To address these problems, researchers developed different strategies such as identifying core node sets (Guo et al.
+2022) and designing scoring functions with a bias towards nodes with large degrees (Luo et al.
 2018).
-In this study, we explore an alternative approach inspired by the the success of large language models (LLMs) in many fields (Gu et al.
+((a))
+((b))
+((c))
+Figure 1: Some shortcomings faced by seed expansion algorithms. Nodes in the same community are marked with the same color.
+In this study, inspired by the success of Large Language Models (LLMs) in many fields (Gu et al.
 2024; Zong and Krishnamachari 2024; Yang et al.
-2022) such as theory-of-mind reasoning (Adhikari et al.
-2020; Ammanabrolu and Riedl 2021) and graphical reasoning (Wang et al.
-2023).
-As far as we know, LLMs have not been applied to local community detection.
-Therefore, we explore the application of large models for local community detection and propose an algorithm, called ComGPT (GPT-guided Local Community Detection).
-This algorithm leverages GPT to guide the detection process, thereby bridging the current gap in LLMs applicability to this field.
-ComGPT integrates the seed expansion algorithm with GPT, leveraging their inherent compatibility: GPT through step-by-step guidance, while the seed expansion method incrementally adds nodes to the community, making their integration highly synergistic.
-However, the application of LLMs to local community detection faces the following challenges: 1) Lack of graph encoding methods for providing LLMs with sufficient community-related graph information.
-To address this, we enhance graph encoding by incorporating two key aspects: graph topology and community information.
-2) Lacks domain knowledge for LLMs in dealing with local community detection.
+2022)(Adhikari et al.
+2020; Ammanabrolu and Riedl 2021)(Wang et al.
+2023), we explore an alternative approach based on LLMs, called ComGPT (GPT-guided seed expansion algorithm).
+ComGPT leverages LLMs to guide the local community detection process to alleviate the seed-dependent problem, community diffusion, and free rider effect.
+The one-question-one-answer modality of LLMs aligns well with the incremental nature of seed expansion algorithms.
+It bridges the current gap in LLMs applicability to the community detection domain.
+Specifically, we leverage LLMs to select nodes to join the community instead of the scoring function commonly used in seed expansion algorithms.
+However, the application of LLMs to select nodes faces the following challenges: 1) Lack of graph encoding methods for providing LLMs with sufficient community-related graph information.
+To address this, we improve the graph encoding method, called Incident (Fatemi, Halcrow, and Perozzi 2023), by incorporating community information.
+2) Lack of domain knowledge for LLMs in dealing with local community detection.
 It limits the capabilities of LLMs.
-To address this, we have designed prompt to assist LLMs in understanding community definitions and how to select nodes.
-3) Lacks efficient interaction between GPT and community detection methods.
-The mode for integrating community detection algorithms with GPT, as well as strategies for reducing the number of input tokens during their interaction, remain unexplored.
-Based on the step-by-step characteristics of GPT and seed expansion algorithm, we design a step-by-step interactive mode.
-To minimize costs by reducing the number of tokens, the input to GPT is not the entire graph but a partial graph that includes the detected community and some of its neighbors.
-Our contribution is summarized as follows:
-  • We propose a GPT-guided local community detection algorithm, called ComGPT.
-ComGPT repeatedly selects potential nodes from the neighbors of the detected community and then uses GPT to identify the node that optimally integrates into the detected community from these selected potential nodes.
-To the best of our knowledge, this is the first work to apply LLMs to the domain of local community detection.
-  • For GPT to understand and detect communities, we design graph encoding method, NSG prompt, and interaction mode.
-The graph encoding method incorporates both graph topology and community information, which enhances GPT to understand the current state of the detected community.
-NSG prompt that integrates community features facilitates GPT’s comprehension of domain knowledge.
-Additionally, an interactive mode is established between GPT and the local community detection method.
+To address this, we design prompts to assist LLMs in understanding community definitions and how to select nodes.
+Additionally, LLMs receive only a partial network as input, not the entire network, to minimize costs by reducing the number of tokens processed.
+Our contributions are summarized as follows:
+  • We propose a GPT-guided seed expansion algorithm called ComGPT, which iteratively uses GPT-3.5-turbo to add nodes to the community.
+It includes four main steps: potential node identification, graph encoding, node selection by GPT-3.5-turbo, and node supplementation by GPT-3.5-turbo.
+To the best of our knowledge, this is the first work to apply LLMs to the domain of community detection.
+  • We design an improved version of the typical graph encoding method, named Incident (Fatemi, Halcrow, and Perozzi 2023), and NSG (Node Selection Guide) prompt.
+This improved version incorporates both graph topology and community information, which enhances LLMs to understand the current state of the detected community.
+NSG prompt that integrates community features facilitates LLMs’s comprehension of domain knowledge.
   • The results on real datasets indicate that ComGPT outperforms comparison methods.
-Ablation study confirm that GPT plays a crucial role in the ComGPT ’s performance.
-Furthermore, prompt comparisons experiments reveal that the designed prompt effectively aids GPT in understanding the graph structure and community situation.
+The ablation study experiments confirm that LLMs play a crucial role in the ComGPT’s performance.
+Furthermore, prompt comparison experiments reveal that the designed prompt effectively aids LLMs in understanding the graph structure and community situation.
 Related work
@@
 2024c, 2023).
-A variety of methods like seed expansion algorithms are developed for detecting local communities (Luo, Wang, and Promislow 2008; Cui et al.
-2014), which expand communities by selecting nodes with the highest quality function values.
-For example, the R method (Clauset 2005) and M method (Luo, Wang, and Promislow 2008) choose nodes based on the highest local modularity R and M, respectively.
-The seed expansion algorithms often faces the seed-dependent problem (Kloumann and Kleinberg 2014), community diffusion, and free rider effect (Sozio and Gionis 2010; Lang and Andersen 2007; Tsourakakis et al.
+A variety of methods, like $k$-core decomposition (Cui et al.
+2014) and seed expansion algorithms (Clauset 2005), are developed for detecting local communities.
+Seed expansion algorithms expand communities by selecting nodes with high scoring function values (Moradi, Olovsson, and Tsigas 2014).
+For example, the R method (Clauset 2005) and M method (Luo, Wang, and Promislow 2008) choose nodes based on the local modularity R and M, respectively.
+The seed expansion algorithms often face the seed-dependent problem (Kloumann and Kleinberg 2014), community diffusion, and free rider effect (Sozio and Gionis 2010; Lang and Andersen 2007; Tsourakakis et al.
 2013).
 To address the seed-dependent problem, researchers developed different strategies such as identifying core node sets (Guo et al.
-2022), finding the nearest nodes with greater centrality (Luo et al.
-2020), and designing membership functions with a bias towards nodes with large degrees (Luo et al.
+2022; Ji, Guo, and Yu 2022), finding the nearest nodes with greater centrality (Luo et al.
+2020), alternating strategy of strong fusion and weak fusion (Shang et al.
+2023), and designing membership functions with a bias towards nodes with large degrees (Luo et al.
 2018).
-Our work mitigates the aforementioned problems by utilizing GPT, distinguishing it from existing studies.
-LLMs
-In recent years, researchers have explored LLMs for node classification (Chen et al.
-2023) and graph reasoning (Zhang 2023).
+Our work mitigates the aforementioned problems by utilizing LLMs, distinguishing them from existing studies.
+LLMs in graph reasoning
+In recent years, researchers have explored the application of LLMs in graph reasoning (Zhang 2023; Chen et al.
+2023).
 Wang et al.
@@
 Furthermore, Guo et al.
-emphasized the design and improvement of graph encoding techniques, which critically affect GPT’s interpretation of input data and its subsequent output (Guo et al.
+emphasized the design and improvement of graph encoding techniques, which critically affect LLMs’ interpretation of input data and its subsequent output (Guo et al.
 2023).
@@
 Existing graph coding methods primarily represent graph topology, which is often insufficient for specific tasks, such as community detection.
-Our work introduces a novel graph encoding method tailored for the community detection domain, which incorporates community structure considerations, distinguishing it from existing approaches.
+Our work introduces a novel graph encoding method based on Incident (Fatemi, Halcrow, and Perozzi 2023), tailored for the community detection domain, that incorporates community structure and distinguishes itself from existing approaches.
 Furthermore, various prompts are designed for specific tasks to make LLMs better adapt to these tasks (Sahoo et al.
-2024), such as the general COT prompt (Wei et al.
-2022) and the BAG prompt (Wang et al.
+2024), such as the general CoT (Chain-of-Thought) prompt (Wei et al.
+2022) and the BaG (Build a Graph) prompt (Wang et al.
 2023) for graph reasoning tasks.
 However, these prompts do not incorporate the domain knowledge required for local community detection and are unsuitable for detecting community structures within networks.
-This paper designs a prompt, named NSG, suitable for the field of local community detection, which is different from the above work.
-Proposed algorithm
-Shortcomings
-Existing heuristic algorithms suffer from the seed-dependent problem (Kloumann and Kleinberg 2014), community diffusion, and free rider effect (Sozio and Gionis 2010) (Lang and Andersen 2007) (Tsourakakis et al.
-2013), which are explained as following.
-(a) Seed-dependent problem
-(b) Community diffusion
-(c) Free rider effect
-Figure 1: Existing shortcomings. Nodes in the same community are marked with the same color.
-Shortcoming 1: Seed-dependent problem (Kloumann and Kleinberg 2014).
-It refers to the quality of the community detected being highly dependent on the initial seed nodes, as illustrated by nodes 5 and 1 in Figure 1(1(a)).
-The communities detected by the M method (Luo, Wang, and Promislow 2008) with nodes 5 and 1 as the starting seed nodes are [5, 11, 6, 7, 17] and [1, 12, 13, 4, 8, 2, 14, 3, 22, 18, 20], respectively.
-The former includes some nodes of the real community and misses many nodes, while the latter is more accurate.
-Shortcoming 2: Community diffusion.
-The community diffusion problem means that the detected community contains nodes from different communities.
-Take node 1 in Figure 1(1(b)) as an example to illustrate the community diffusion problem.
-The community detected by the M method with node 1 as the starting seed node is [1, 2, 3, 4, 7, 8, 9, 13, 14, 15, 16, 17, 18, 20], which includes nodes from two different communities.
-Shortcoming 3: Free rider effect (Sozio and Gionis 2010) (Lang and Andersen 2007) (Tsourakakis et al. 2013).
-The free rider effect refers to mistakenly adding unrelated communities to enhance the fitness of a detected community.
-We use the community derived from seed node 7, as shown in Figure 1(1(c)), to illustrate it.
-The target community consists of nodes 5, 6, 7, 8, and 9.
-However, to achieve a higher fitness value, the detected community not only contains nodes in the target community [5, 6, 7, 8, 9] but also erroneously includes all nodes in the unrelated community [1, 2, 3, 4, 10, 11].
-ComGPT
+This paper designs a prompt, named NSG, suitable for the field of local community detection, which is different from the above prompts.
+Approach
 Figure 2: The diagram of ComGPT.
-We propose a local community detection algorithm based on GPT, called ComGPT.
-Figure 2 shows the main steps of ComGPT.
-It includes four main steps: 1) potential nodes identification; 2) graph encoding; 3) node selection by GPT; 4) node supplementation by GPT.
-Potential nodes identification aims to pinpoint potential nodes from the neighbors of the current community.
-Graph encoding involves transforming the graph structure into a textual format interpretable by GPT.
-Node selection is that GPT selects the node that optimally integrates into the community from potential nodes.
-Node supplement involves GPT selecting which nodes to add to the community when no potential nodes remain.
+It includes four main steps: 1) Potential node identification aims to pinpoint potential nodes from the neighbors of the current community.
+2) Graph encoding involves transforming the graph structure into a textual format interpretable by GPT-3.5-turbo.
+3) Node selection by GPT-3.5-turbo is that GPT-3.5-turbo selects the node that optimally integrates into the community from potential nodes.
+4) Node supplementation by GPT-3.5-turbo involves GPT-3.5-turbo selecting which nodes to add to the community when no potential nodes remain.
+We introduce a GPT-guided seed expansion algorithm named ComGPT, based on GPT-3.5-turbo. This section begins with an overview of ComGPT, followed by detailed descriptions of each step of ComGPT.
+Overview
 Algorithm 1 ComGPT
-Input: Network $G$, Seed node $v_{s}$
+Input: Network $G$, Seed node $v_{s}$, Prompts NSG and NSU for node selection and supplementation, respectively, Questions Qsel and Qsup for node selection and supplementation, respectively
 Output: Community $C$
@@
 4:     while $PN\neq\emptyset$ do
-5:        $gtxt\leftarrow$ graph encoding with $C$ and $N$
-6:        $v\leftarrow$ node selection by GPT from $PN$
+5:        $gtxt\leftarrow$ graph encoding with ($C\cup PN$, $C$, $PN$)
+6:        $v\leftarrow$ node selection by GPT-3.5-turbo with ($gtxt$, NSG, Qsel )
 7:        add $v$ to $C$
@@
 9:     add $C$ to $cands$
-10:     $v\leftarrow$ node supplementation by GPT
-11:     if $v$ is not null then
-12:        add $v$ to $C$
-13:     else
-14:        break
-15:     end if
-16:  end while
-17:  $C\leftarrow$ community with highest $M$ in $cands$
-18:  return  $C$
-Algorithm 1 shows the overall process of ComGPT.
-For convenience, we briefly introduce the notations used in this paper.
-Let $G$ denote the network, $v_{s}$ denote the seed node, $C$ denote the detected community, $v$ denote the node returned by the node selection or node appending steps, $gtxt$ denote the text of the graph structure obtained through graph encoding, $PN$ denote the set of potential nodes, $cands$ store the community each time $PN$ is empty.
-Initially, community $C$ consists of only one seed node $v_{s}$ and the $cands$ is empty (line 1).
-First, ComGPT selects potential nodes $PN$ from neighbors of current community $C$ for GPT (line 3).
-The local network, i.e., the subgraph consisting of the nodes of the current community $C$ and $PN$, is converted into text, i.e., $gtxt$ (line 5).
-Utilizing $gtxt$ and NSG (Node Select Guide) prompt, GPT selects the node $v$ from potential nodes $PN$ and adds it to the current community $C$ (line 6,7).
+10:     $N\leftarrow$ first and second-order neighbors of $C$
+11:     $PN\leftarrow$ nodes selected by modularity $M$
+12:     $gtxt\leftarrow$ graph encoding with ($C\cup N$, $C$, $PN$)
+13:     $v\leftarrow$ node supplementation by GPT-3.5-turbo with ($gtxt$, NSU, Qsup)
+14:     if $v$ is not null then
+15:        add $v$ to $C$
+16:     else
+17:        break
+18:     end if
+19:  end while
+20:  $C\leftarrow$ community with highest $M$ in $cands$
+21:  return  $C$
+As shown in Figure 2, ComGPT contains four main steps.
+To better understand the interactions among these steps, Algorithm 1 shows the overall process of ComGPT.
+Initially, community $C$ contains seed node $v_{s}$ and $cands$ is empty (line 1), where $cands$ stores the community whenever the potential node identification step returns an empty set.
+First, ComGPT selects potential nodes $PN$ with the step of potential node identification (line 3), used for subsequent graph encoding.
+In order for GPT-3.5-turbo to understand the graph structure and community information, the local network, i.e., the subgraph consisting of the nodes in both current community $C$ and $PN$ as well as edges between these nodes, is converted into text, i.e., $gtxt$ (line 5).
+Utilizing $gtxt$, NSG (Node Select Guide) prompt, and Qsel, GPT-3.5-turbo selects node $v$ from potential nodes $PN$ and adds it to the current community $C$ (lines 6-7).
 The above steps are repeated until $PN$ is empty.
-After the potential nodes $PN$ is empty, the community $C$ is added to the candidate communities $cands$ (line 9).
-When $PN$ is empty, it does not necessarily imply that the neighboring nodes of the community are weakly connected to the community.
-Instead, it could be that the nodes have relatively high degrees, which marginally reduces the $M$ value.
-To address this scenario, ComGPT identifies nodes through node supplementation (line 10).
-If GPT identifies suitable node $v$ to add to the community, then ComGPT adds $v$ to $C$ and continues to expand current community $C$ (line 12).
-Otherwise, ComGPT stops community expansion (line 14).
-After the above process is completed, ComGPT selects the community with the highest $M$ value from $cands$ as the final community (line 17).
+If $PN$ is empty, the community $C$ is added to the $cands$ (line 9).
+The fact that $PN$ is empty does not necessarily imply that the neighboring nodes of the community are weakly connected to the community.
+Therefore, based on $gtxt$, NSU (Node SUpplementatio) prompt, and Qsup, ComGPT continues to identify nodes through node supplementation (lines 10-13).
+If GPT-3.5-turbo identifies suitable node $v$ to add to the community, then ComGPT adds $v$ to $C$ and continues to expand current community $C$ (line 15).
+Otherwise, GPT-3.5-turbo returns “null” and ComGPT stops community expansion (line 17).
+After the above process is completed, ComGPT selects the community with the highest local modularity $M$ calculated by formula (1) from $cands$ as the final community (line 20).
 Potential nodes identification
-To mitigate the high computational cost associated with processing nodes through GPT, ComGPT identifies potential nodes that may belong to the current community and provides these nodes to GPT for selection.
-Different fitness functions are designed in the seed expansion algorithm to select nodes, which can be used to choose potential nodes, such as modularity $M$ (Luo, Wang, and Promislow 2008), $R$ (Clauset 2005), and $Q_{lcd}$ (Shang et al.
+To mitigate the cost associated with processing nodes through GPT-3.5-turbo, ComGPT identifies potential nodes that may belong to the current community and provides these nodes to GPT-3.5-turbo for selection.
+Different scoring functions are designed in the seed expansion algorithms to select nodes (Moradi, Olovsson, and Tsigas 2014), which can be used to choose potential nodes, such as modularity $M$ (Luo, Wang, and Promislow 2008), $R$ (Clauset 2005), and $Q_{lcd}$ (Shang et al.
 2023).
-ComGPT uses local modularity $M_{C}$ to select potential nodes, which is defined as the ratio of the number of internal edges in community $C$ to the number of external edges of community $C$.
-Nodes that are satisfied $\Delta M=M_{C\cup\{v\}}-M_{C}>0$ are regarded as potential nodes.
-During the initial expansion of the community, since nodes with high degrees often struggle to achieve $\Delta M>0$, these nodes are not considered potential nodes $PN$.
-To alleviate this preference, these nodes are added to potential nodes based on their closeness value to the current community $C$.
-The closeness value $f_{v}$ between node $v$ and the community $C$ is calculated as
-$f_{v}=\frac{|C\cap N(v)|}{|C|},$ (1)
-$|C|$ represents the number of nodes in the community and $N(v)$ represents the neighbors of node $v$.
-The value of $f_{v}$ ranges from 0 to 1.
-Higher values indicate a greater proportion of nodes connected to those within the community.
-The initial potential nodes consist of nodes with $\Delta M>0$ and nodes with high $f_{v}$ values.
-To further reduce the number of nodes processed by GPT, $k$ nodes are selected from them to form a potential node’s set.
+ComGPT uses local modularity $M_{C}$ (Luo, Wang, and Promislow 2008) to select potential nodes.
+$M_{C}$ is calculated as,
+$M_{C}=\frac{e_{ic}}{e_{oc}},$ (1)
+where $e_{ic}$ represents the number of internal edges in community $C$ and $e_{oc}$ represents the number of external edges of community $C$.
+For each first-order neighbor node $v$ of community $C$, if node $v$ satisfies $\Delta M=M_{C\cup\{v\}}-M_{C}>0$, then it is considered a potential node.
+To further reduce the number of nodes processed by GPT-3.5-turbo, $k$ nodes with the largest $M$ are selected from the above potential nodes to form a potential node’s set $PN$.
 Graph encoding
-Figure 3: Graph encoding.
-The purpose of graph encoding is to convert graph structure data and information about the community $C$ into text that is easily understandable by GPT.
+Figure 3: An toy example of graph encoding.
+The middle part illustrates the text corresponding to graph topological structure.
+The lower part illustrates the text corresponding to supplementary community information.
+The purpose of graph encoding is to convert graph structure data and information about the community $C$ into text that is easily understandable by GPT-3.5-turbo.
 The graph data includes two aspects.
 1) Graph topological structure.
-We adopt the Incident method (Fatemi, Halcrow, and Perozzi 2023) to let GPT understand the graph’s topological structure.
-The Incident method represents the graph topology by introducing the first-order neighbors of each node.
-To reduce computational overhead, the graph structure inputted to GPT is a local network, i.e., the subgraph consisting of the nodes of the current community $C$ and $PN$.
-For clarity, the middle part of Figure 3 illustrates the specific graph encoding: “G describes a graph among nodes:1, 2, 3, 5, 6, 7.
-Node 1 is connected to nodes 2,3”.
-To reduce the number of tokens, some conjunctions that do not affect the understanding of GPT are removed.
+We adopt Incident (Fatemi, Halcrow, and Perozzi 2023) to let GPT-3.5-turbo understand the graph’s topological structure.
+Incident represents the graph topology by introducing the first-order neighbors of each node.
+To reduce computational overhead, the graph structure input to GPT-3.5-turbo is a local network.
+For the node selection by GPT-3.5-turbo step, the local network comprises nodes in $C\cup PN$ and the edges between them.
+For the node supplementation by GPT-3.5-turbo step, the local network consists of the nodes in $C\cup N$ and edges between these nodes, where $N$ includes first and second-order neighbors of community $C$.
 2) Supplementary knowledge.
-We describe node information from a community perspective to enhance GPT’s understanding of community structure.
-This involves outlining which nodes are within the community, those that are outside, and detailing their specific external connections.
-For clarity, the lower part of Figure 3 illustrates graph encoding for community $C$ = [1, 2, 3] and potential nodes $PN$ = [5, 6, 7]: “Nodes in the current community:[1,2,3].
-The outside nodes contain:[5,6,7].
-Node 5 is connected to nodes in community:[1,3]”.
-Node selection by GPT
-Node selection refers to choosing the most suitable node for community expansion, based on the prompt and graph data.
-Below, we introduce node selection through prompts that teach GPT to accurately select nodes.
-1. The more a potential node is connected to other potential nodes, the higher the likelihood of its selection: This prompt could alleviate the seed-dependent problem and the free rider effect.
+We describe node information from a community perspective to enhance GPT-3.5-turbo’s understanding of community structure.
+This step involves outlining which nodes are within the community, those that are outside, and the connections of outside nodes to both within and outside the community.
+The outside nodes are potential nodes in $PN$ from the step of potential nodes identification.
+The text corresponding to the topological structure and the text corresponding to the supplementary information are concatenated to form $gtxt$.
+To better understanding, an example is given in Figure 3.
+Node selection by GPT-3.5-turbo
+Node selection refers to choosing the most suitable node for community expansion, based on NSG prompt, graph text $gtxt$, and question Qsel.
+Graph text $gtxt$ is obtained by the step of graph encoding with the local network, $C$, and $PN$, where the local network consists of the nodes in $C\cup PN$ and edges between these nodes.
+Below, we introduce NSG prompt and question Qsel.
+  • Prompt 1: The more an outside node is connected to other outside nodes, the higher the likelihood of its selection.
+The outside nodes provided to GPT-3.5-turbo are potential nodes from the step of potential nodes identification.
+This prompt could alleviate the seed-dependent problem and the free rider effect.
     The reason for the seed-dependent problem is low-degree node preference.
-Take seed node 5 in Figure 1(1(a)) and M method (Luo, Wang, and Promislow 2008) as an example to illustrate it.
+Take seed node 5 in Figure 1(a) and M method (Luo, Wang, and Promislow 2008) as an example to illustrate it.
 Nodes 1, 7, and 11 are each connected to the node in the community [5] by an edge, and node 11, having the smallest degree, is selected so that the community has the fewest external edges.
-Thus, the M method selects node 11 to join the community.
+Thus, the M method (Luo, Wang, and Promislow 2008) selects node 11 to join the community.
 Nodes that connect to many other potential nodes are usually high-degree nodes, which help to avoid bias towards low-degree nodes and thereby alleviate the seed-dependent problem.
-Taking Figure 1(1(a)) as an example, the community includes node 5, with nodes 1, 7, and 11 regarded as potential nodes.
+Nodes 1, 7, and 11 are potential nodes for community [5].
 According to prompt 1, node 1 will be selected over the mistakenly considered node 11.
-The reason for the free rider effect is the neglect of the connectivity of neighbor nodes of the detected community.
-Nodes that connect to many other neighbor nodes are consideration of their association with other potential nodes, which alleviates the free rider effect.
-Taking Figure 1(1(c)) as an example, the community contains node 7, with nodes 4, 6, and 8 regarded as potential nodes.
+    The reason for the free rider effect is adding nodes that are sparsely connected to the community to the community.
+Nodes that are connected to multiple potential nodes indirectly maintain close connections to the community, which alleviates the free rider effect.
+Taking Figure 1(c) as an example, the community contains node 7, with nodes 4, 6, and 8 regarded as potential nodes.
 According to prompt 1, nodes 6 and 8 have a higher probability of being selected.
 Node 4 is a node that may cause the free rider effect, and will not be selected.
-2.
-Prioritize selecting nodes that are connected to multiple nodes within the community: This prompt could alleviate the community diffusion problem.
+  • Prompt 2: Prioritize selecting outside nodes that are connected to multiple nodes within the community.
+The outside nodes mentioned in prompts are potential nodes from the step of potential nodes identification.
+This prompt could alleviate the community diffusion problem.
 The reason for the community diffusion is the insufficient consideration of the connections between nodes to be added to the community and the nodes within the community.
 Therefore, the prompt considers the connection between potential nodes and the community, thereby alleviating the community diffusion problem.
-Taking Figure 1(1(b)) as an example, the community includes nodes 1 and 3, with nodes 2, 5, 7, 13, 15, and 18 regarded as potential nodes, where node 13 or 18 may lead to community diffusion.
+Taking Figure 1(b) as an example, the community includes nodes 1 and 3, with nodes 2, 5, 7, 13, 15, and 18 regarded as potential nodes, where nodes 13 or 18 may lead to community diffusion.
 According to prompt 2, nodes 2 and 5 have a higher probability of being selected, rather than mistakenly selecting nodes 13 or 18.
-The first prompt is to consider the closeness with neighboring nodes outside the community, and the second prompt is to consider the closeness with nodes within the community.
-This is consistent with the characteristics of the community being close inside and sparse outside, thus helping GPT understand the characteristics of the nodes to be selected.
-Node supplementation by GPT
-When $PN$ is empty, ComGPT uses GPT to supplement nodes.
+    The first prompt is to consider the closeness of the node to potential nodes outside the community, and the second prompt is to consider the closeness of the node to nodes within the community.
+This is consistent with the characteristics of the community being close inside, thus helping GPT-3.5-turbo understand the characteristics of the nodes to be selected.
+To guide GPT-3.5-turbo to provide relevant answers, the question Qsel for selecting nodes is designed: “You’re doing local community detection.
+Based on the graph data and prompt, please select a node that you think is most likely to belong to the current community $C$ for community expansion.
+Provide a detailed explanation.”.
+Based on the above prompts, the text $gtxt$ obtained by graph encoding method, and the question Qsel, GPT selects the node $v$ from potential nodes.
+Node supplementation by GPT-3.5-turbo
 Node supplementation involves evaluating whether to continue the expansion process or terminate it.
-If continuation is warranted, it also determines which additional nodes should be selected.
-In terms of graph encoding, the graph encoding method mentioned above, including graph topological structure and supplementary knowledge, is used.
-The difference is that to provide GPT with sufficient information to make a decision, the input graph text corresponds to graph topology that includes the current community and its first and second-order neighbors.
-The nodes to be selected by GPT are the three nodes with the largest M values among the first-order neighbor nodes of the community.
-In terms of prompt, it is based on the definition of the community and whether these nodes can maintain the cohesion of the community.
-If GPT determines that any of these three nodes warrant inclusion in the community, it returns the node’s number; otherwise, it returns NULL.
+If continuing, it also determines which additional nodes should be selected.
+Next, we introduce the information needed for graph encoding, NSU prompt, and question Qsup in turn.
+For graph encoding, to provide GPT-3.5-turbo with sufficient information to make a decision, the local network consists of the nodes in $C\cup N$ and edges between these nodes, where $N$ includes first and second-order neighbors of community $C$.
+Besides, ComGPT forcibly adds the $k$ nodes with the highest $M$ values from the first neighbors of community $C$ to $PN$.
+NSU prompt is based on the definition of the community and whether these nodes can maintain the cohesion of the community.
+To guide GPT-3.5-turbo to provide relevant answers, the question Qsel for supplementing nodes is designed: “Please analyze whether these nodes should be added to the community $C$.
+The probability of not adding nodes is higher.
+But it doesn’t mean you always refuse to add nodes.
+If you think there is a suitable node, please output its node number.”.
+Based on the above prompts, the text obtained by graph encoding method, and the question Qsel, if GPT-3.5-turbo determines that any of nodes warrant inclusion in the community, it returns the node’s number; otherwise, it returns “null”.
 Experiments
-To evaluate the performance of our algorithm, we conducted some experiments on real datasets.
-We first outline the experimental settings, which include datasets, comparison algorithms, and evaluation metrics.
-We then present the experimental results and provide corresponding analyses.
 Experimental settings
-dataset M R DMF_M DMF_R LCDNN LCDMD ComGPT
-Dolphins 0.4685 0.4505 0.7255 0.7242 0.7558 0.7235 0.7885
-Polbooks 0.5257 0.5030 0.7295 0.7286 0.7463 0.7301 0.7586
-Football 0.7146 0.7107 0.8881 0.8893 0.8948 0.8134 0.8406
-Amazon 0.7706 0.7554 0.8353 0.8361 0.8372 - 0.8385
-DBLP 0.4185 0.4185 0.4032 0.4009 0.3152 - 0.4247
-Table 1: Comparisons with other methods in terms of fscore.
+ Dataset M R DMF_M DMF_R LCDNN LS ComGPT
+Dolphins Jaccard 0.3370 0.3128 0.6270 0.6320 0.6438* 0.7735 0.7035
+Fscore 0.4673 0.4478 0.7254 0.7241 0.7558* 0.8385 0.7684
+Polbooks Jaccard 0.4266 0.3986 0.6449 0.6437 0.6673 0.6919 0.6468*
+Fscore 0.5184 0.4960 0.7294 0.7285 0.7463 0.7764 0.7320*
+Football Jaccard 0.6117 0.6134 0.8546 0.8579 0.7547 0.2406 0.7571*
+Fscore 0.6706 0.6699 0.8881 0.8893 0.8255* 0.3593 0.8103
+Amazon Jaccard 0.6961 0.6771 0.7633 0.7686 0.7665* 0.6506 0.7734
+Fscore 0.7706 0.7554 0.8353 0.8361* 0.8372 0.7420 0.8421
+DBLP Jaccard 0.3239 0.3207* 0.2945 0.2928 0.2208 0.0650 0.3259
+Fscore 0.4185 0.4185 0.4032 0.4009 0.3152 0.0850 0.4359
+Table 1: Experimental results on five datasets.
+On each dataset, the highest Fscore (Jaccard) is highlighted in bold, the second highest one is highlighted by underline, and the third highest one is highlighted with a superscript asterisk.
 Datasets
@@
 The Football network contains 115 nodes and 613 edges.
-Nodes within this network represent college football teams, communities correspond to different leagues, and edges correspond to games played between them.
+Nodes represent college football teams, communities correspond to different leagues, and edges correspond to games played between them.
 The Polbooks network is a co-purchasing network for books.
@@
 An edge between two books indicates that they are often purchased together by customers.
-It contains 105 nodes and 441 edges.The Amazon network represents product co-purchasing relationships, consisting of 334,863 nodes and 925,872 edges.
-In this network, nodes representing products and edges indicate which products are frequently co-purchased.
+It contains 105 nodes and 441 edges.
+The Amazon network represents product co-purchasing relationships, consisting of 334,863 nodes and 925,872 edges.
+Nodes representing products and edges indicate which products are frequently co-purchased.
 The DBLP network is a co-authorship network comprising 317,080 nodes and 1,049,866 edges.
-Here, nodes represent authors, and edges indicate that two authors have co-authored a paper.
-Comparison algorithms
-To evaluate the performance of ComGPT, we compare it with five local community detection algorithms: M method (Luo, Wang, and Promislow 2008), R method (Clauset 2005), DMF_M (Luo et al.
+Here, nodes represent authors and edges indicate that two authors have co-authored a paper.
+Baselines
+To evaluate the performance of ComGPT, we compare it with M method (Luo, Wang, and Promislow 2008), R method (Clauset 2005), DMF_M (Luo et al.
 2018), DMF_R (Luo et al.
 2018), LCDNN (Luo et al.
-2020), and LCDMD (Guo et al.
-2022).
-ComGPT uses gpt-3.5-turbo-0125 and sets $k_{f}$ to 0.5.
-Due to monetary costs of GPT-3.5-turbo-0125, we execute ComGPT twice for each starting seed node to handle GPT hallucinations.
+2020), and LS (Shi et al.
+2024).
+Implementation Details
+ComGPT uses gpt-3.5-turbo-0125.
+The number of potential nodes $k$ is set to 10.
+All experiments are conducted under Windows 10 operating system (CPU: 11th Gen Intel(R) Core(TM) i7-11700K @ 3.60GHZ, Memory: 32 GB.).
+Due to the monetary costs of GPT-3.5-turbo-0125, we execute ComGPT twice for each starting seed node to handle GPT hallucinations.
 Evaluation metrics
-The evaluation metrics precision, recall, fscore (Powers 2020) are used to evaluate the detected communities.
+The evaluation metrics Fscore (Powers 2011) and Jaccard (Jaccard 1912) are used to evaluate the detected communities.
 For an overlapped seed node, all the communities it belongs to are integrated as a real local community.
 Fscore is the harmonic mean of precision and recall.
-The values of precision, recall, and fscore are between 0 and 1.
+The values of Fscore and Jaccard are between 0 and 1.
 The larger value implies a better algorithm performance.
@@
 Results on networks
-For the Dolphins, Polbooks, and Football datasets, each node is selected as a seed node separately.
-For the Amazon and DBLP datasets, We randomly select 200 nodes from the top 5,000 communities.
-The average fscore of all seed nodes is calculated.
+For the Dolphins, Polbooks, and Football datasets, all node are selected as seed nodes.
+For the Amazon and DBLP datasets, we randomly select 200 nodes from the top 5,000 communities as seed nodes.
+The average Fscore and Jaccard of all seed nodes is calculated.
 The detailed experimental results are shown in Table 1.
-Due to the unavailability of the LCDMD (Guo et al.
-2022) source code, results for the Dolphins, Polbooks, and Football datasets are cited from (Guo et al.
-2022), in which results for the Amazon and DBLP datasets are not available.
-Table 1 shows that ComGPT outperforms the M method, R method, DMF_M, DMF_R, LCDNN, and LCDMD on most datasets.
-ComGPT is superior to other methods because GPT has excellent comprehension to find the suitable node based on the NSG prompt.
+Table 1 shows that ComGPT outperforms the M method, R method, DMF_M, DMF_R, LCDNN, and LS on most datasets.
+Among all comparison methods, ComGPT exhibits the best performance on the Amazon and DBLP datasets and ranks second or third on the other three datasets.
+Although LS achieves the best performance on the Dolphins and Polbooks datasets, its performance on the remaining three datasets is the lowest among all comparison methods.
+Overall, ComGPT is superior to LS.
+ComGPT is superior to other methods because GPT-3.5-turbo has excellent comprehension to find the suitable node based on the NSG prompt.
 On the Football dataset, ComGPT performs worse than DMF_M, DMF_R, and LCDNN.
-It is because Football contains several tightly interconnected communities, which leads GPT to erroneously identify them as a single community.
-Except for the DBLP dataset, DMF_M, DMF_R, LCDNN, and LCDMD perform better than the M method and R method.
-The reason why DMF_M, DMF_R, LCDNN, and LCDMD outperform the M method and R method on most datasets is that the latter two methods maximize local modularity (M and R), making them sensitive to the choice of seed nodes, known as seed-dependent problem.
-In contrast, DMF_M, DMF_R, LCDNN, and LCDMD implement various strategies to alleviate this problem.
-Specifically, DMF_M, and DMF_R, utilize dynamic membership functions to select nodes, which tend to favor nodes with large degrees.
-LCDNN expands the community based on the fuzzy relationships between nodes and NGC nodes.
-Meanwhile, LCDMD introduces a new modularity density to effectively mine the core parts of a community.
+It is because Football contains several tightly interconnected communities, which leads GPT-3.5-turbo to erroneously identify them as a single community.
+Except for the DBLP dataset, DMF_M, DMF_R, and LCDNN perform better than the M method and R method.
+The reason why DMF_M, DMF_R, and LCDNN outperform the M method and R method on most datasets is that the latter two methods maximize local modularity ($M$ and $R$), making them sensitive to the choice of seed nodes, known as seed-dependent problem.
+In contrast, DMF_M, DMF_R, and LCDNN implement various strategies to alleviate this problem.
+Specifically, DMF_M, and DMF_R, utilize membership functions to select nodes with large degrees.
+LCDNN expands the community based on the fuzzy relations between nodes and NGC nodes.
 Ablation study
-Method GPT SK NSG precision recall fscore
-1 NO NO NO 0.6824 0.7246 0.6978
-2 YES NO YES 0.7571 0.8277 0.7832
-3 YES YES NO 0.5911 0.7026 0.6221
-ComGPT YES YES YES 0.8440 0.8520 0.8445
-Table 2: Results of ablation experiment on the football dataset.
-To explore the necessity of GPT, SK (Supplementary knowledge in graph coding), and NSG prompt in ComGPT, we conduct ablation experiments on the Football dataset with 50 randomly selected seed nodes.
-Three comparative models are constructed as follows: 1) Method 1 does not use GPT.
-The first node in $PN$ will be selected for community expansion until $PN$ is empty; 2) Method 2 only maintains GPT and NSG prompt, without SK; and 3) Method 3 only maintains GPT and SK, without NSG prompt.
-When calculating precision and recall, it is necessary to obtain the intersection between the detected community, which may include duplicate nodes, and the true community.
-In this calculation, duplicate nodes in the detected community are regarded as a single node.
+Method GPT SK NSG NSU Fscore Jaccard
+1 NO NO NO NO 0.4666 0.3352
+2 YES NO YES YES 0.3711 0.2414
+3 YES YES NO YES 0.4855 0.3299
+4 YES YES YES NO 0.7134 0.6307
+ComGPT YES YES YES YES 0.7688 0.7009
+Table 2: Results of ablation experiment on the Dolphins dataset.
+To explore the necessity of GPT-3.5-turbo, SK (Supplementary knowledge in graph coding), NSG, and NSU prompts in ComGPT, we conduct ablation experiments on the Dolphins dataset with 50 randomly selected seed nodes.
+Four simplified versions of ComGPT are constructed as follows: 1) Method 1 without GPT-3.5-turbo.
+The first node in $PN$ is selected for community expansion until $PN$ is empty; 2) Method 2 without SK; 3) Method 3 without NSG prompt; and 4) Method 4 without NSU prompt.
+When calculating Fscore and Jaccard, it is necessary to obtain the intersection between the detected community, which may include duplicate nodes, and the true community.
+Here, duplicate nodes in the detected community are regarded as a single node.
 To prevent cyclic selection of duplicate nodes, community expansion is terminated once the number of duplicate nodes exceeds one-third of the number of nodes in the community.
-Table 2 shows the experimental results of Methods 1, 2, 3, and ComGPT.
-Overall, ComGPT outperforms Methods 1, 2, and 3, indicating that GPT, SK, and NSG prompt play a crucial role in ComGPT, respectively.
-ComGPT is better than Method 1 in terms of precision, recall, and fscore, which demonstrates the effectiveness and importance of GPT.
+Table 2 shows the experimental results of Methods 1, 2, 3, 4, and ComGPT.
+Overall, ComGPT outperforms Methods 1, 2, 3, and 4, indicating that GPT-3.5-turbo, SK, NSG, and NSU prompt play a crucial role in ComGPT, respectively.
+ComGPT is better than Method 1 in terms of Fscore and Jaccard, which demonstrates the effectiveness and importance of GPT-3.5-turbo.
 The reason ComGPT outperforms Method 2 is that its graph coding incorporates community knowledge.
-Compared with graph encoding methods that contain only the graph topology, supplementing community knowledge enables the GPT to better understand the community structure and thus make more accurate judgments.
-ComGPT also outperforms Method 3, suggesting that NSG prompt can significantly enhance its performance in local community detection.
+Compared with graph encoding methods that contain only the graph topology, supplementing community knowledge enables the GPT-3.5-turbo to better understand the community structure and thus make more accurate judgments.
+ComGPT also outperforms Method 3 and 4, suggesting that NSG and NSU prompts can enhance its performance in local community detection.
 Prompt comparisons
- Method Dolphins Polbooks Football
-Zero-shot P 0.7195 0.5793 0.5911
-R 0.3911 0.2330 0.7026
-F 0.4690 0.3261 0.6221
-Few-shot P 0.7136 0.5658 0.6592
-R 0.3840 0.3358 0.7600
-F 0.4556 0.4070 0.6894
-COT P 0.7483 0.5924 0.6120
-R 0.4397 0.2633 0.7205
-F 0.5173 0.3572 0.6512
-BAG P 0.7242 0.6005 0.6530
-R 0.4704 0.2338 0.7589
-F 0.5192 0.3315 0.6870
-NSG P 0.9592 0.7788 0.8440
-R 0.7356 0.7559 0.8520
-F 0.7964 0.7506 0.8445
-Table 3: Results of different prompts. P, R, and F are the abbreviations of recall, precision, and fscore, respectively.
+Prompts Description
+Few-shot It is given a few demonstrations of the task
+at inference time as conditioning.
+Zero-shot No demonstrations are allowed and only an instruction in natural language is given to the model.
+CoT Let’s think step by step.
+BaG Let’s construct a graph with the nodes and edges first.
+Table 3: Description of the prompts.
+ Prompts Dolphins Polbooks Football
+Zero-shot Jaccard 0.3299 0.2191 0.5829
+Fscore 0.4855 0.3490 0.6774
+Few-shot Jaccard 0.3584 0.2703 0.7145
+Fscore 0.5128 0.4078 0.7764
+CoT Jaccard 0.3250 0.2078 0.6168
+Fscore 0.4785 0.3347 0.7196
+BaG Jaccard 0.2772 0.1813 0.6050
+Fscore 0.4235 0.2988 0.7131
+NSG Jaccard 0.7009 0.6624 0.8045
+Fscore 0.7688 0.7418 0.8555
+Table 4: Results of different prompts.
 To test the performance of the NSG prompt within ComGPT, we replaced the NSG prompt with four typical prompt approaches: Zero-shot (Brown et al.
@@
 2022), and BaG (Wang et al.
-2023) prompts.
+2023) prompts, as shown in Table 3.
 These alternate versions of the algorithm serve as baseline models for comparison.
 We conducted comparative experiments by randomly selecting 50 seed nodes in the Polbooks, Dolphins, and Football datasets.
-The detected communities may contain duplicate nodes, and the indicator values are calculated in the same way as in the ablation study.
-Table 3 shows that the performance of NSG is better than that of Zero-shot, Few-shot, COT, and BAG prompts.
-Specifically, on Polbooks and Football datasets, the Few-shot outperforms the Zero-shot.
-Few-shot is worse than Zero-shot on the Dolphins dataset.
-This is because each dataset has a different structure.
-If the provided examples match the community structure characteristics of the dataset, the performance of GPT is improved.
-Conversely, if they do not match, GPT’s performance is diminished.
-The reason COT and BAG outperform Zero-shot on most datasets is that COT enhances GPT’s reasoning ability by guiding it to think step-by-step and BAG helps GPT better digest graph information.
-On Dolphins, Polbooks, and Football datasets, NSG outperforms other prompts.
+The detected communities may contain duplicate nodes, and the metrics values are calculated in the same way as in the ablation study.
+Table 4 shows that the performance of ComGPT with NSG is better than that of ComGPT with Zero-shot, Few-shot, CoT, and BaG prompts.
+Specifically, on Football dataset, the ComGPT with CoT and BaG outperform the ComGPT with Zero-shot.
+The ComGPT with CoT and BaG is worse than ComGPT with Zero-shot on Dolphins and Polbooks datasets.
+Advanced prompts may have limited performance improvement or even counterproductive effects on complex graph reasoning tasks.
+For instance, advanced prompts like CoT increase the number of reasoning steps for GPT-3.5-turbo, which potentially raises the likelihood of errors in the final reasoning conclusion due to mistakes in one of the intermediate steps.
+The reason ComGPT with Few-shot outperforms ComGPT with Zero-shot on most datasets is that we provided some examples of how the task should be executed.
+These examples guide the model by showing the correct process needed to solve the task.
+On Dolphins, Polbooks, and Football datasets, ComGPT with NSG outperforms ComGPT with other prompts.
 This is because the design of CoT and BaG is oriented towards universal graph reasoning problems so that no relevant guidance or knowledge is utilized in dealing with local community detection.
@@
 As the $k$ increases, the performance of ComGPT remains generally stable, though with some fluctuations.
-Fluctuations are relatively low across different values, except for $k$ = 13, which exhibits higher variability due to the randomness of GPT output.
-To confirm this, we conduct repeated experiments for $k$ = 13 and the results are nearly identical to those obtained with other values of $k$.
+Fluctuations are relatively low across different values, except for $k$ = 10, which exhibits higher variability due to the randomness of GPT-3.5-turbo output.
+To confirm this, we conduct repeated experiments for $k$ = 10 and the results are nearly identical to those obtained with other values of $k$.
 The above analysis suggests that the performance of ComGPT is less affected by the parameter $k$.
 Conclusion
-We explored the potential of using GPT to assist in local community detection and developed a GPT-guided methodology, termed ComGPT.
+We explored using GPT-3.5-turbo to assist in seed expansion algorithm and developed a GPT-guided methodology, termed ComGPT.
 Additionally, a graph encoding method suitable for local community detection is explored, which enhances the existing graph encoding by incorporating community knowledge.
 Furthermore, the NSG prompt is designed to enhance LLMs’ understanding of community domain knowledge.
-Experimental results demonstrate that ComGPT outperforms the comparison algorithms.
-Moreover, the designed graph encoding and prompt help GPT understand the domain knowledge of graph topology and community discovery, respectively.
+Experimental results confirm that the designed graph encoding method and prompts help GPT-3.5-turbo understand the domain knowledge and detect local community structure, respectively.

### v3

@@
 Abstract
-Large Language Models (LLMs), like GPT, have demonstrated the ability to understand graph structures and have achieved excellent performance in various graph reasoning tasks, such as node classification.
+Large Language Models (LLMs), like GPT-3.5-turbo, have demonstrated the ability to understand graph structures and have achieved excellent performance in various graph reasoning tasks, such as node classification.
 Despite their strong abilities in graph reasoning tasks, they lack specific domain knowledge and have a weaker understanding of community-related graph information, which hinders their capabilities in the community detection task.
-Moreover, local community detection algorithms based on seed expansion, referred to as seed expansion algorithms, often face the seed-dependent problem, community diffusion, and free rider effect.
+Moreover, local community detection algorithms based on seed expansion, referred to as seed expansion algorithms, often face several shortcomings, including the seed-dependent problem, community diffusion, and free rider effect.
 To use LLMs to overcome the above shortcomings, we explore a GPT-guided seed expansion algorithm named ComGPT.
-ComGPT iteratively selects potential nodes by local modularity $M$ from the detected community’s neighbors, and subsequently employs LLMs to choose the node to join the detected community from these selected potential nodes.
-To address the above issues faced by LLMs, we improve graph encoding method, called Incident, by incorporating community knowledge to improve LLMs’s understanding of community-related graph information.
-Additionally, we design the NSG (Node Selection Guide) prompt to enhance LLMs’ understanding of community characteristics.
-Experimental results demonstrate that ComGPT outperforms the comparison methods, thereby confirming the effectiveness of the improved graph encoding method and prompts.
-Introduction
-Identifying community structures from networks has wide applications, such as identifying protein complexes (Yin et al.
-2017).
-Local community detection has gained attention due to its ability to quickly identify the community containing a specific node (Ni et al.
-2024a, b).
-Various methods are designed for detecting local community, including seed expansion algorithms (Luo, Wang, and Promislow 2008), heat diffusion techniques (Kloster and Gleich 2014), and non-negative matrix factorization algorithms (Wang et al.
-2011).
+ComGPT iteratively selects potential nodes by local modularity from the detected community’s neighbors, and subsequently employs LLMs to choose the node from these selected potential nodes to join the detected community.
+To improve LLMs’ understanding of community-related graph information, we propose ComIncident, a graph encoding method that incorporates community knowledge and is designed for the community detection task.
+Additionally, we design the Node Selection Guide (NSG) prompt to enhance LLMs’ understanding of community characteristics.
+Experimental results demonstrate that ComGPT outperforms the baselines, thereby confirming the effectiveness of the ComIncident and the NSG prompt.
+I Introduction
+Identifying community structures from networks has wide applications, such as identifying protein complexes [1].
+Local community detection has gained attention due to its ability to quickly identify the community containing a specific node [2, 3].
+Various methods are designed for detecting local community, including seed expansion algorithms [4], pagerank based techniques [5], flow-propagation based techniques [6, 7], and non-negative matrix factorization algorithms [8].
 Among these methods, the seed expansion algorithms have garnered widespread acclaim for their versatility and ease of implementation.
-The seed expansion algorithms typically enlarge the community by selecting a node with the high scoring function value at each step (Moradi, Olovsson, and Tsigas 2014).
-However, seed expansion algorithms often face the seed-dependent problem (Kloumann and Kleinberg 2014), community diffusion, and free rider effect (Sozio and Gionis 2010; Lang and Andersen 2007; Tsourakakis et al.
-2013), shown in Figure 1.
-The seed-dependent problem refers to the quality of the community detected being highly dependent on the initial seed nodes, as illustrated by nodes 5 and 1 in Figure 1(a).
-The community diffusion problem means that the detected community contains nodes from different communities, as illustrated by node 1 in Figure 1(b).
-The free rider effect refers to mistakenly adding unrelated communities to enhance the scoring function value of a detected community, as illustrated by node 7 in Figure 1(c).
-To address these problems, researchers developed different strategies such as identifying core node sets (Guo et al.
-2022) and designing scoring functions with a bias towards nodes with large degrees (Luo et al.
-2018).
-((a))
-((b))
-((c))
-Figure 1: Some shortcomings faced by seed expansion algorithms. Nodes in the same community are marked with the same color.
-In this study, inspired by the success of Large Language Models (LLMs) in many fields (Gu et al.
-2024; Zong and Krishnamachari 2024; Yang et al.
-2022)(Adhikari et al.
-2020; Ammanabrolu and Riedl 2021)(Wang et al.
-2023), we explore an alternative approach based on LLMs, called ComGPT (GPT-guided seed expansion algorithm).
-ComGPT leverages LLMs to guide the local community detection process to alleviate the seed-dependent problem, community diffusion, and free rider effect.
-The one-question-one-answer modality of LLMs aligns well with the incremental nature of seed expansion algorithms.
-It bridges the current gap in LLMs applicability to the community detection domain.
-Specifically, we leverage LLMs to select nodes to join the community instead of the scoring function commonly used in seed expansion algorithms.
-However, the application of LLMs to select nodes faces the following challenges: 1) Lack of graph encoding methods for providing LLMs with sufficient community-related graph information.
-To address this, we improve the graph encoding method, called Incident (Fatemi, Halcrow, and Perozzi 2023), by incorporating community information.
-2) Lack of domain knowledge for LLMs in dealing with local community detection.
+It typically enlarges the community by selecting a node with the high scoring function value at each step [9].
+Figure 1: Seed-dependent problem.
+The community detected by M method [4] with starting nodes 5 is {5, 11, 6, 7, 17}.
+This community misses many nodes.
+Nodes in the same community are marked with the same color.
+Figure 2: Community diffusion.
+The community detected by the M method [4] with node 1 as the starting seed node is {1, 2, 3, 4, 7, 8, 9, 13, 14, 15, 16, 17, 18, 20}, which incorrectly includes nodes from two different communities.
+Nodes in the same community are marked with the same color.
+Figure 3: Free rider effect.
+The target community consists of nodes 5, 6, 7, 8, and 9.
+The detected community not only contains nodes in the target community {5, 6, 7, 8, 9} but also erroneously includes all nodes in the unrelated community {1, 2, 3, 4, 10, 11}.
+Nodes in the same community are marked with the same color.
+However, seed expansion algorithms often face the seed-dependent problem [10], community diffusion, and free rider effect [11, 12, 13], which are shown in Figure 1, 2, and 3, respectively.
+The seed-dependent problem refers to the quality of the community detected being highly dependent on the initial seed nodes, as illustrated by node 5 in Figure 1.
+The community diffusion problem means that the detected community contains nodes from different communities, as illustrated by node 1 in Figure 2.
+The free rider effect refers to mistakenly adding unrelated communities to enhance the scoring function value of a detected community, as illustrated by node 7 in Figure 3.
+To address these problems, researchers developed different strategies such as identifying core node sets [14], finding the nearest nodes with greater centrality [15], alternating strategy of strong fusion and weak fusion [16], and so on.
+The above studies rely on predefined heuristic rules, which may fail to adapt to diverse structures across networks.
+Recently, Large Language Models (LLMs) have emerged as powerful reasoning engines [17, 18, 19, 20, 21, 22].
+Unlike these heuristic-based approaches, LLMs do not rely on a fixed set of rules and can be prompted to reason about diverse network structures.
+Therefore, we explore a GPT-guided seed expansion approach based on LLMs, called ComGPT.
+ComGPT leverages LLMs to guide the detection of a local community.
+The one-question-one-answer modality of LLMs aligns well with the incremental nature of seed expansion algorithms, making it naturally suitable for this task.
+This approach bridges the current gap in LLMs applicability to the community detection domain.
+Specifically, we leverage LLMs to select nodes to join the community instead of the scoring function used in seed expansion algorithms.
+However, the application of LLMs to select nodes faces the following challenges:
+  1. 1.
+    Lack of graph encoding methods for providing LLMs with sufficient community-related graph information.
+To address this, we propose ComIncident (Community-Enriched Incident), a variant of the graph encoding method Incident [23], which incorporates community information.
+  2. 2.
+    Lack of domain knowledge for LLMs in dealing with local community detection.
 It limits the capabilities of LLMs.
 To address this, we design prompts to assist LLMs in understanding community definitions and how to select nodes.
-Additionally, LLMs receive only a partial network as input, not the entire network, to minimize costs by reducing the number of tokens processed.
+  3. 3.
+    Token count constraints. To address this, LLMs receive only a partial network as input, not the entire network, to minimize costs by reducing the number of tokens processed.
 Our contributions are summarized as follows:
+  1. 1.
     We propose a GPT-guided seed expansion algorithm called ComGPT, which iteratively uses GPT-3.5-turbo to add nodes to the community.
@@
 To the best of our knowledge, this is the first work to apply LLMs to the domain of community detection.
-  • We design an improved version of the typical graph encoding method, named Incident (Fatemi, Halcrow, and Perozzi 2023), and NSG (Node Selection Guide) prompt.
-This improved version incorporates both graph topology and community information, which enhances LLMs to understand the current state of the detected community.
-NSG prompt that integrates community features facilitates LLMs’s comprehension of domain knowledge.
-  • The results on real datasets indicate that ComGPT outperforms comparison methods.
+  2. 2.
+    We design an graph encoding method, named ComIncident, and NSG (Node Selection Guide) prompt.
+ComIncident is the first graph encoding method designed for the community detection task, which enhances LLMs to understand the current state of the detected community.
+The NSG prompt, which integrates community features, facilitates LLMs’s comprehension of domain knowledge.
+  3. 3.
+    The results on real datasets indicate that ComGPT outperforms baselines.
 The ablation study experiments confirm that LLMs play a crucial role in the ComGPT’s performance.
 Furthermore, prompt comparison experiments reveal that the designed prompt effectively aids LLMs in understanding the graph structure and community situation.
-Related work
-Local community detection
-Local community detection focuses on discovering the community to which the seed node belongs (Ni et al.
-2024c, 2023).
-A variety of methods, like $k$-core decomposition (Cui et al.
-2014) and seed expansion algorithms (Clauset 2005), are developed for detecting local communities.
-Seed expansion algorithms expand communities by selecting nodes with high scoring function values (Moradi, Olovsson, and Tsigas 2014).
-For example, the R method (Clauset 2005) and M method (Luo, Wang, and Promislow 2008) choose nodes based on the local modularity R and M, respectively.
-The seed expansion algorithms often face the seed-dependent problem (Kloumann and Kleinberg 2014), community diffusion, and free rider effect (Sozio and Gionis 2010; Lang and Andersen 2007; Tsourakakis et al.
-2013).
-To address the seed-dependent problem, researchers developed different strategies such as identifying core node sets (Guo et al.
-2022; Ji, Guo, and Yu 2022), finding the nearest nodes with greater centrality (Luo et al.
-2020), alternating strategy of strong fusion and weak fusion (Shang et al.
-2023), and designing membership functions with a bias towards nodes with large degrees (Luo et al.
-2018).
-Our work mitigates the aforementioned problems by utilizing LLMs, distinguishing them from existing studies.
-LLMs in graph reasoning
-In recent years, researchers have explored the application of LLMs in graph reasoning (Zhang 2023; Chen et al.
-2023).
-Wang et al.
-demonstrated LLMs’ graph reasoning abilities and designed the NLGraph dataset (Wang et al.
-2023) to evaluate these capabilities.
-Furthermore, Guo et al.
-emphasized the design and improvement of graph encoding techniques, which critically affect LLMs’ interpretation of input data and its subsequent output (Guo et al.
-2023).
-Therefore, Adjacency and Incident (Fatemi, Halcrow, and Perozzi 2023) are designed to encode graph structured data into text for LLMs.
-Adjacency uses integer node encoding and parenthesis edge encoding.
-Incident uses integer node encoding and incident edge encoding.
-Existing graph coding methods primarily represent graph topology, which is often insufficient for specific tasks, such as community detection.
-Our work introduces a novel graph encoding method based on Incident (Fatemi, Halcrow, and Perozzi 2023), tailored for the community detection domain, that incorporates community structure and distinguishes itself from existing approaches.
-Furthermore, various prompts are designed for specific tasks to make LLMs better adapt to these tasks (Sahoo et al.
-2024), such as the general CoT (Chain-of-Thought) prompt (Wei et al.
-2022) and the BaG (Build a Graph) prompt (Wang et al.
-2023) for graph reasoning tasks.
-However, these prompts do not incorporate the domain knowledge required for local community detection and are unsuitable for detecting community structures within networks.
-This paper designs a prompt, named NSG, suitable for the field of local community detection, which is different from the above prompts.
-Approach
+The rest of the paper is organized as follows.
+Section II presents the proposed algorithm.
+Section III reports experimental results.
+Section IV summarizes the related works.
+Section V concludes the paper.
+II Approach
 Figure 4: The diagram of ComGPT.
@@
 4) Node supplementation by GPT-3.5-turbo involves GPT-3.5-turbo selecting which nodes to add to the community when no potential nodes remain.
-We introduce a GPT-guided seed expansion algorithm named ComGPT, based on GPT-3.5-turbo. This section begins with an overview of ComGPT, followed by detailed descriptions of each step of ComGPT.
-Overview
+We introduce a GPT-guided seed expansion algorithm named ComGPT.
+This section begins with an overview of ComGPT, followed by detailed descriptions of each step of ComGPT.
+Table I lists the important notations used in the paper.
+II-A Overview
+TABLE I: Important notations
+Notation Definition
+$G$ network
+$C$ community
+NN neighbor node of $C$
+$v_{s}$ seed node
+PNPN potential node
+candscands communities each time $PN$ is empty
+gtxtgtxt graph text obtained by the graph encoding step
+NSGNSG node selection guide prompt
+NSUNSU node supplementation prompt
+$Q_{sel}$ node selection question
+$Q_{sup}$ node supplementation question
+As shown in Figure 4, ComGPT contains four main steps: potential node identification, graph encoding, node selection by GPT-3.5-turbo, and node supplementation by GPT-3.5-turbo.
+The idea of ComGPT uses potential node identification step (Section II-B) to select potential nodes from the neighbors of the current community.
+Then, it uses node selection by GPT-3.5-turbo step (Section II-D) and node supplementation by GPT-3.5-turbo step (Section II-E) selects nodes to join the community.
+Since GPT-3.5-turbo cannot directly process graph structures and community structures, the graph encoding step (Section II-C) is used to encode the graph structure and community structure into text that GPT-3.5-turbo can understand.
+Algorithm 1 shows the overall process of ComGPT.
+Initially, community $C$ contains seed node $v_{s}$ and candscands is empty (line 1), where candscands stores the communities each time $PN$ is empty.
+First, ComGPT selects potential nodes $PN$ with the step of potential node identification (line 3), used for subsequent graph encoding.
+In order for GPT-3.5-turbo to understand the graph structure and community information, the local network, i.e., the subgraph consisting of the nodes in both current community $C$ and $PN$ as well as edges between these nodes, is converted into text, i.e., gtxtgtxt (line 5).
+Utilizing gtxtgtxt, GPT-3.5-turbo selects node $v$ from potential nodes $PN$ and adds it to the current community $C$ (lines 6-7).
+The above steps are repeated until $PN$ is empty.
+Then the community $C$ is added to the candscands (line 9).
+$PN$ being empty does not necessarily imply that the neighboring nodes of the community are weakly connected to it.
+Therefore, ComGPT continues to identify nodes through node supplementation (lines 10-13).
+If GPT-3.5-turbo identifies suitable node $v$, ComGPT adds $v$ to $C$ and continues to expand current community $C$ (line 15).
+Otherwise, GPT-3.5-turbo returns “null” and ComGPT stops community expansion (line 17).
+After the above process is completed, ComGPT selects the community with the highest local modularity $M$ calculated by formula (1) from candscands as the final community (line 20).
+During the expansion process, if no new node is selected in potential nodes identification step (Section II-B), i.e., $PN$ is empty (line 3), ComGPT jumps out of the inner while loop.
+Thereafter, if the node selected in the node supplementation by GPT-3.5-turbo is null (line 14), ComGPT terminates.
 Algorithm 1 ComGPT
@@
 Output: Community $C$
-1:  $C\leftarrow{\{v_{s}\}}$, $cands\leftarrow\emptyset$
+[…]
 2:while True do
-3:     $PN\leftarrow$ potential nodes identification with $C$
+3:  $PN\leftarrow$ potential nodes identification with $C$ // Section II-B
 4:  while $PN\neq\emptyset$ do
-5:        $gtxt\leftarrow$ graph encoding with ($C\cup PN$, $C$, $PN$)
-6:        $v\leftarrow$ node selection by GPT-3.5-turbo with ($gtxt$, NSG, Qsel )
+5:   $gtxt\leftarrow$ graph encoding with ($\cup PN, CC$, PNPN) // Section II-C
+6:   $v\leftarrow$ node selection by GPT-3.5-turbo with (gtxtgtxt, NSG, Qsel ) // Section II-D
 7:   add $v$ to $C$
 8:  end while
-9:     add $C$ to $cands$
+9:  add $C$ to candscands
 10:  $N\leftarrow$ first and second-order neighbors of $C$
 11:  $PN\leftarrow$ nodes selected by modularity $M$
-12:     $gtxt\leftarrow$ graph encoding with ($C\cup N$, $C$, $PN$)
-13:     $v\leftarrow$ node supplementation by GPT-3.5-turbo with ($gtxt$, NSU, Qsup)
+12:  $gtxt\leftarrow$ graph encoding with ($\cup N, CC$, PNPN) // Section II-C
+13:  $v\leftarrow$ node supplementation by GPT-3.5-turbo with (gtxtgtxt, NSU, Qsup) // Section II-E
 14:  if $v$ is not null then
@@
 19:end while
-20:  $C\leftarrow$ community with highest $M$ in $cands$
+[…]
 21:return $C$
-As shown in Figure 2, ComGPT contains four main steps.
-To better understand the interactions among these steps, Algorithm 1 shows the overall process of ComGPT.
-Initially, community $C$ contains seed node $v_{s}$ and $cands$ is empty (line 1), where $cands$ stores the community whenever the potential node identification step returns an empty set.
-First, ComGPT selects potential nodes $PN$ with the step of potential node identification (line 3), used for subsequent graph encoding.
-In order for GPT-3.5-turbo to understand the graph structure and community information, the local network, i.e., the subgraph consisting of the nodes in both current community $C$ and $PN$ as well as edges between these nodes, is converted into text, i.e., $gtxt$ (line 5).
-Utilizing $gtxt$, NSG (Node Select Guide) prompt, and Qsel, GPT-3.5-turbo selects node $v$ from potential nodes $PN$ and adds it to the current community $C$ (lines 6-7).
-The above steps are repeated until $PN$ is empty.
-If $PN$ is empty, the community $C$ is added to the $cands$ (line 9).
-The fact that $PN$ is empty does not necessarily imply that the neighboring nodes of the community are weakly connected to the community.
-Therefore, based on $gtxt$, NSU (Node SUpplementatio) prompt, and Qsup, ComGPT continues to identify nodes through node supplementation (lines 10-13).
-If GPT-3.5-turbo identifies suitable node $v$ to add to the community, then ComGPT adds $v$ to $C$ and continues to expand current community $C$ (line 15).
-Otherwise, GPT-3.5-turbo returns “null” and ComGPT stops community expansion (line 17).
-After the above process is completed, ComGPT selects the community with the highest local modularity $M$ calculated by formula (1) from $cands$ as the final community (line 20).
-Potential nodes identification
-To mitigate the cost associated with processing nodes through GPT-3.5-turbo, ComGPT identifies potential nodes that may belong to the current community and provides these nodes to GPT-3.5-turbo for selection.
-Different scoring functions are designed in the seed expansion algorithms to select nodes (Moradi, Olovsson, and Tsigas 2014), which can be used to choose potential nodes, such as modularity $M$ (Luo, Wang, and Promislow 2008), $R$ (Clauset 2005), and $Q_{lcd}$ (Shang et al.
-2023).
-ComGPT uses local modularity $M_{C}$ (Luo, Wang, and Promislow 2008) to select potential nodes.
+II-B Potential nodes identification
+To mitigate the cost associated with processing nodes through GPT-3.5-turbo, ComGPT identifies potential nodes $PN$ that may belong to the current community and provides $PN$ to GPT-3.5-turbo for selection.
+Different scoring functions are designed in the seed expansion algorithms to select nodes [9], which can be used to choose potential nodes, such as modularity $M$ [4], $R$ [24], and $Q_{lcd}$ [16].
+ComGPT uses local modularity $M_{C}$ [4] to select potential nodes.
 $M_{C}$ is calculated as,
@@
 To further reduce the number of nodes processed by GPT-3.5-turbo, $k$ nodes with the largest $M$ are selected from the above potential nodes to form a potential node’s set $PN$.
-Graph encoding
-Figure 3: An toy example of graph encoding.
-The middle part illustrates the text corresponding to graph topological structure.
-The lower part illustrates the text corresponding to supplementary community information.
+II-C Graph encoding
+Figure 5: An toy example of graph encoding. The middle part illustrates the text corresponding to topology graph. The lower part illustrates the text corresponding to supplementary knowledge.
 The purpose of graph encoding is to convert graph structure data and information about the community $C$ into text that is easily understandable by GPT-3.5-turbo.
-The graph data includes two aspects.
-1) Graph topological structure.
-We adopt Incident (Fatemi, Halcrow, and Perozzi 2023) to let GPT-3.5-turbo understand the graph’s topological structure.
-Incident represents the graph topology by introducing the first-order neighbors of each node.
-To reduce computational overhead, the graph structure input to GPT-3.5-turbo is a local network.
-For the node selection by GPT-3.5-turbo step, the local network comprises nodes in $C\cup PN$ and the edges between them.
-For the node supplementation by GPT-3.5-turbo step, the local network consists of the nodes in $C\cup N$ and edges between these nodes, where $N$ includes first and second-order neighbors of community $C$.
+We design graph encoding method, called ComIncident, for the community detection task.
+ComIncident includes two aspects: 1) Topology graph.
+The graph topology is represented by introducing the first-order neighbors of each node, as in Incident [23].
 2) Supplementary knowledge.
 We describe node information from a community perspective to enhance GPT-3.5-turbo’s understanding of community structure.
-This step involves outlining which nodes are within the community, those that are outside, and the connections of outside nodes to both within and outside the community.
-The outside nodes are potential nodes in $PN$ from the step of potential nodes identification.
-The text corresponding to the topological structure and the text corresponding to the supplementary information are concatenated to form $gtxt$.
-To better understanding, an example is given in Figure 3.
-Node selection by GPT-3.5-turbo
-Node selection refers to choosing the most suitable node for community expansion, based on NSG prompt, graph text $gtxt$, and question Qsel.
-Graph text $gtxt$ is obtained by the step of graph encoding with the local network, $C$, and $PN$, where the local network consists of the nodes in $C\cup PN$ and edges between these nodes.
-Below, we introduce NSG prompt and question Qsel.
-  • Prompt 1: The more an outside node is connected to other outside nodes, the higher the likelihood of its selection.
-The outside nodes provided to GPT-3.5-turbo are potential nodes from the step of potential nodes identification.
+It involves outlining which nodes are within the community, those that are outside, and the connections of potential nodes to both within and outside the community.
+The topological structure of local network and the supplementary information are merged to form gtxtgtxt.
+ComIncident is an improved version of Incident.
+Incident utilizes only the topology graph, whereas ComIncident incorporates both the topology graph and supplementary knowledge.
+To better understand ComIncident, an example is given in Figure 5.
+For the node selection by GPT-3.5-turbo step (Section II-D), the local network comprises nodes in $C\cup PN$ and the edges between them.
+For the node supplementation by GPT-3.5-turbo step (Section II-E), the local network consists of nodes in $C\cup N$ and edges between these nodes, where $N$ includes the first and second-order neighbors of community $C$.
+II-D Node selection by GPT-3.5-turbo
+Node selection refers to choosing the most suitable node for community expansion, based on NSG (Node Selection Guide) prompt, graph text gtxtgtxt, and question Qsel.
+Graph text gtxtgtxt is obtained by the graph encoding step (Section II-C) with the local network, $C$, and $PN$, where the local network consists of the nodes in $C\cup PN$ and edges between these nodes.
+Below, we introduce the NSG prompt and question Qsel, where the NSG prompt contains NSG_1 and NSG_2.
+Prompt NSG_1: The more an outside node is connected to other outside nodes, the higher the likelihood of its selection.
+The outside nodes provided to GPT-3.5-turbo are potential nodes from the potential nodes identification step (Section II-B).
 This prompt could alleviate the seed-dependent problem and the free rider effect.
 The reason for the seed-dependent problem is low-degree node preference.
-Take seed node 5 in Figure 1(a) and M method (Luo, Wang, and Promislow 2008) as an example to illustrate it.
-Nodes 1, 7, and 11 are each connected to the node in the community [5] by an edge, and node 11, having the smallest degree, is selected so that the community has the fewest external edges.
-Thus, the M method (Luo, Wang, and Promislow 2008) selects node 11 to join the community.
-Nodes that connect to many other potential nodes are usually high-degree nodes, which help to avoid bias towards low-degree nodes and thereby alleviate the seed-dependent problem.
-Nodes 1, 7, and 11 are potential nodes for community [5].
-According to prompt 1, node 1 will be selected over the mistakenly considered node 11.
+Take seed node 5 in Figure 1 and M method [4] as an example to illustrate it.
+Nodes 1, 7, and 11 are connected to nodes within the community [5].
+The M method [4] selects node 11, having the smallest degree, to join the community, ensuring the fewest external edges of the community.
+With the incorporation of the prompt NSG_1, nodes that connect to many other potential nodes are prioritized.
+Consequently, node 1 is selected over the mistakenly considered node 11.
+That is, prompt NSG_1 helps to avoid bias towards low-degree nodes and thereby alleviates the seed-dependent problem.
 The reason for the free rider effect is adding nodes that are sparsely connected to the community to the community.
 Nodes that are connected to multiple potential nodes indirectly maintain close connections to the community, which alleviates the free rider effect.
-Taking Figure 1(c) as an example, the community contains node 7, with nodes 4, 6, and 8 regarded as potential nodes.
-According to prompt 1, nodes 6 and 8 have a higher probability of being selected.
+Taking Figure 3 as an example, the community contains node 7, with nodes 4, 6, and 8 regarded as potential nodes.
+According to NSG_1, nodes 6 and 8 have a higher probability of being selected.
 Node 4 is a node that may cause the free rider effect, and will not be selected.
-  • Prompt 2: Prioritize selecting outside nodes that are connected to multiple nodes within the community.
-The outside nodes mentioned in prompts are potential nodes from the step of potential nodes identification.
-This prompt could alleviate the community diffusion problem.
+Prompt NSG_2: Prioritize selecting outside nodes that are connected to multiple nodes within the community.
+The outside nodes mentioned in prompts are potential nodes from the potential nodes identification step (Section II-B).
+This prompt could alleviate the community diffusion.
 The reason for the community diffusion is the insufficient consideration of the connections between nodes to be added to the community and the nodes within the community.
 Therefore, the prompt considers the connection between potential nodes and the community, thereby alleviating the community diffusion problem.
-Taking Figure 1(b) as an example, the community includes nodes 1 and 3, with nodes 2, 5, 7, 13, 15, and 18 regarded as potential nodes, where nodes 13 or 18 may lead to community diffusion.
-According to prompt 2, nodes 2 and 5 have a higher probability of being selected, rather than mistakenly selecting nodes 13 or 18.
-    The first prompt is to consider the closeness of the node to potential nodes outside the community, and the second prompt is to consider the closeness of the node to nodes within the community.
-This is consistent with the characteristics of the community being close inside, thus helping GPT-3.5-turbo understand the characteristics of the nodes to be selected.
-To guide GPT-3.5-turbo to provide relevant answers, the question Qsel for selecting nodes is designed: “You’re doing local community detection.
+Taking Figure 2 as an example, the community includes nodes 1 and 3, with nodes 2, 5, 7, 13, 15, and 18 regarded as potential nodes, where nodes 13 or 18 may lead to community diffusion.
+According to NSG_2, nodes 2 and 5 are more likely selected, rather than mistakenly selecting nodes 13 or 18.
+The NSG_1 considers the closeness of the node and potential nodes that are highly likely to be added into the community as ComGPT proceeds.
+The NSG_2 focuses on the closeness of the node and nodes within the community.
+These two prompts align with the characteristics of communities where nodes in the community are tightly connected, thus helping GPT-3.5-turbo understand the characteristics of the nodes to be selected.
+To guide GPT-3.5-turbo to provide relevant answers, the question Qsel is designed:
+You’re doing local community detection.
 Based on the graph data and prompt, please select a node that you think is most likely to belong to the current community $C$ for community expansion.
 Provide a detailed explanation.”.
-Based on the above prompts, the text $gtxt$ obtained by graph encoding method, and the question Qsel, GPT selects the node $v$ from potential nodes.
-Node supplementation by GPT-3.5-turbo
+Based on the above prompts, the graph text gtxtgtxt (Section II-C), and the question Qsel, GPT-3.5-turbo selects the node $v$ from potential nodes.
+II-E Node supplementation by GPT-3.5-turbo
 Node supplementation involves evaluating whether to continue the expansion process or terminate it.
 If continuing, it also determines which additional nodes should be selected.
-Next, we introduce the information needed for graph encoding, NSU prompt, and question Qsup in turn.
-For graph encoding, to provide GPT-3.5-turbo with sufficient information to make a decision, the local network consists of the nodes in $C\cup N$ and edges between these nodes, where $N$ includes first and second-order neighbors of community $C$.
-Besides, ComGPT forcibly adds the $k$ nodes with the highest $M$ values from the first neighbors of community $C$ to $PN$.
-NSU prompt is based on the definition of the community and whether these nodes can maintain the cohesion of the community.
-To guide GPT-3.5-turbo to provide relevant answers, the question Qsel for supplementing nodes is designed: “Please analyze whether these nodes should be added to the community $C$.
+Next, we introduce NSU (Node SUpplementation) prompt and question Qsup in turn.
+NSU prompt is based on the definition of the community and whether these nodes can maintain the cohesion of the community, as follows:
+Definition of community: nodes in the same community are tightly connected, while the nodes in different communities are sparsely connected.
+The more a node connects with nodes within a community, the more it can increase the community’s cohesion.
+Conversely, the more it connects with nodes outside the community, the more it can decrease the community’s cohesion.
+If you choose to add a node to the community, it should make the connections within the community tighter.”.
+The question Qsel for supplementing nodes is designed as:
+Please analyze whether these nodes should be added to the community $C$.
 The probability of not adding nodes is higher.
@@
 If you think there is a suitable node, please output its node number.”.
-Based on the above prompts, the text obtained by graph encoding method, and the question Qsel, if GPT-3.5-turbo determines that any of nodes warrant inclusion in the community, it returns the node’s number; otherwise, it returns “null”.
-Experiments
-Experimental settings
- Dataset M R DMF_M DMF_R LCDNN LS ComGPT
-Dolphins Jaccard 0.3370 0.3128 0.6270 0.6320 0.6438* 0.7735 0.7035
-Fscore 0.4673 0.4478 0.7254 0.7241 0.7558* 0.8385 0.7684
-Polbooks Jaccard 0.4266 0.3986 0.6449 0.6437 0.6673 0.6919 0.6468*
-Fscore 0.5184 0.4960 0.7294 0.7285 0.7463 0.7764 0.7320*
-Football Jaccard 0.6117 0.6134 0.8546 0.8579 0.7547 0.2406 0.7571*
-Fscore 0.6706 0.6699 0.8881 0.8893 0.8255* 0.3593 0.8103
-Amazon Jaccard 0.6961 0.6771 0.7633 0.7686 0.7665* 0.6506 0.7734
-Fscore 0.7706 0.7554 0.8353 0.8361* 0.8372 0.7420 0.8421
-DBLP Jaccard 0.3239 0.3207* 0.2945 0.2928 0.2208 0.0650 0.3259
-Fscore 0.4185 0.4185 0.4032 0.4009 0.3152 0.0850 0.4359
-Table 1: Experimental results on five datasets.
-On each dataset, the highest Fscore (Jaccard) is highlighted in bold, the second highest one is highlighted by underline, and the third highest one is highlighted with a superscript asterisk.
-Datasets
-Five real-world datasets, including Dolphins (Lusseau 2003), Football (Girvan and Newman 2002), Polbooks (Krebs 2004), Amazon (Yang and Leskovec 2012), and DBLP (Yang and Leskovec 2012) are utilized to test the performance of algorithms.
+Based on the NSU prompt, the text obtained by the graph encoding step (Section II-C), and the question Qsel, GPT-3.5-turbo determines whether any of the nodes warrant inclusion in the community.
+If so, it returns the node’s number; otherwise, it returns ‘null’.
+ComGPT aims to explore the application of LLMs to local community detection.
+Specifically, it skillfully combines LLMs with the seed expansion algorithm.
+Different from traditional seed expansion methods that rely on predefined and rigid heuristic rules, ComGPT leverages the implicit semantic reasoning capability of LLMs to guide the seed expansion process.
+Notably, we do not simply replace a certain rule of the seed expansion algorithm with LLMs.
+Instead, ComGPT provides the LLM with guiding principles rather than explicit, step-by-step rules.
+This leverages the LLM’s inherent reasoning capability to grasp the nuanced semantics of concepts like ‘tightness’ and ‘community membership’.
+Consequently, the LLMs can be prompted to reason about diverse network structures.
+II-F Complexity analysis
+The time complexity analysis is as follows.
+Assuming there are $n$ nodes in the local community and the average degree of nodes is $d$.
+The potential node identification step involves traversing the neighbors of nodes within the local community and calculating modularity M, with a time complexity of O($nd^{2}$).
+The graph encoding step requires generating textual descriptions for the current community and potential nodes, with a time complexity proportional to the number of nodes and their degrees, i.e., O($nd$).
+The core of the node selection and node supplementation steps involves calling the GPT-3.5-turbo API.
+Since there is no literature defining the time complexity of GPT-3.5-turbo API calls, this portion of the time complexity remains unestimated.
+ComGPT repeats the above steps $n$ times to get the final community.
+Excluding the GPT-3.5-turbo API calls, the time complexity of ComGPT is O($n^{2}d^{2}$).
+The space complexity analysis is as follows.
+The potential node identification step stores local community, neighbors of local community, and $k$ potential nodes.
+In ComGPT, $k$ is treated as a constant with a default value of 10; therefore, the space complexity of this step is O($n+nd+k$).
+The graph encoding step requires storing the current community along with its potential nodes and the connections between nodes, resulting in a space complexity of O($nd+kd$).
+The node selection and node supplementation steps only store the response results from GPT-3.5-turbo, with a space complexity of O($1$).
+Therefore, the overall space complexity of ComGPT is O($nd+kd$).
+III Experiments
+In this section, we first introduce the experimental settings, including datasets, baselines, implementation details, and evaluation metrics.
+Then, we give the experimental results and corresponding analyses.
+III-A Experimental settings
+III-A1 Datasets
+TABLE II: Statistics of the datasets.
+Networks Nodes Edges $\bar{d}$ $|C|$ $\mu$ $\bar{Q}$
+Dolphins 62 159 5.13 31 0.0383 0.3734
+Football 115 613 10.66 9.58 0.3638 0.5539
+Polbooks 105 441 8.4 35 0.1812 0.4149
+Amazon 334863 925872 5.53 30.23 0.1216 0.4441
+DBLP 317080 1049866 6.62 53.41 0.2147 0.4453
+Five real-world datasets, including Dolphins [25], Football [26], Polbooks [27], Amazon [28], and DBLP [28] are utilized to test the performance of algorithms.
+Table II shows the detailed information of five real networks, including the number of nodes, the number of edges, the average degree $\bar{d}$, the average community size $|C|$, the mixing parameter $\mu$ [29], and the modularity $Q$ [30].
 The Dolphins network is based on observations of associations between pairs of dolphins, where edges represent associations between dolphins.
-It contains 62 nodes and 159 edges.
-The Football network contains 115 nodes and 613 edges.
+The Football network is based on college football matches.
 Nodes represent college football teams, communities correspond to different leagues, and edges correspond to games played between them.
@@
 An edge between two books indicates that they are often purchased together by customers.
-It contains 105 nodes and 441 edges.
-The Amazon network represents product co-purchasing relationships, consisting of 334,863 nodes and 925,872 edges.
+The Amazon network represents product co-purchasing relationships.
 Nodes representing products and edges indicate which products are frequently co-purchased.
-The DBLP network is a co-authorship network comprising 317,080 nodes and 1,049,866 edges.
+The DBLP network is a co-authorship network.
 Here, nodes represent authors and edges indicate that two authors have co-authored a paper.
-Baselines
-To evaluate the performance of ComGPT, we compare it with M method (Luo, Wang, and Promislow 2008), R method (Clauset 2005), DMF_M (Luo et al.
-2018), DMF_R (Luo et al.
-2018), LCDNN (Luo et al.
-2020), and LS (Shi et al.
-2024).
-Implementation Details
+III-A2 Baselines
+To evaluate the performance of ComGPT, we compare it with M method [4], R method [24], Louvain [31], DMF_M [29], DMF_R [29], Leiden [32], LCDNN [15], LS [33], and FLMIG [34], introduced as follows:
+  1. 1.
+    M method [4] and R method [24].
+The M method and R method are greedy algorithms based on modularity.
+The given node constitutes the initial local community at first.
+M method (R method) iteratively selects and adds the node that maximizes the modularity MM($R$) to the local community until no node will decrease modularity MM($R$).
+  2. 2.
+    Louvain [31].
+The Louvain first treats each node as an individual community and iteratively adjusts node assignments to maximize modularity until stability is reached.
+Then, it aggregates communities into new nodes, constructs a simplified graph, and repeats the optimization process until modularity converges.
+  3. 3.
+    DMF_M and DMF_R[29]. DMF_M and DMF_R analyze the formation process of the local community and divide it into three stages. Each stage corresponds to a different dynamic membership function.
+  4. 4.
+    Leiden [32].
+Leiden is a derivative version of Louvain.
+Unlike Louvain, Leiden ensures well-connected communities through a refinement step.
+It then contracts communities into new nodes, constructs a simplified graph, and repeats the process until convergence, achieving more stable partitions.
+  5. 5.
+    LCDNN [15].
+LCDNN relies on NGC nodes.
+The nodes are added to the local community based on the fuzzy relations between them and their NGC nodes.
+At each time, if the largest fuzzy relation is no less than half of the average fuzzy relation of the current local network, one node with the largest fuzzy relation is added to the local community.
+  6. 6.
+    LS [33].
+LS is built on the notion of local dominance, where low-degree nodes are assigned to the basin of influence of high-degree nodes.
+Utilizing this principle, LS identifies community centers based on node degree and other local information.
+Finally, it divides communities according to the identified community centers.
+  7. 7.
+    FLMIG [34]. FLMIG is also a derivative version of Louvain. FLMIG enhances the Louvain Prune heuristic using an iterated greedy (IG) framework to maximize modularity in non-overlapping communities.
+For baselines that require parameter settings, we select the optimal parameter within the given range as the final value.
+For Louvain, the parameter that controls the size of the communities takes values within [0.3, 0.5, 0.8, 1, 1.5].
+For Leiden, the maximum total size of nodes in a community is tuned within [0, 50, 100, 150, 200].
+For FLMIG, the parameters are set as explicitly provided by the authors.
+III-A3 Implementation Details
 ComGPT uses gpt-3.5-turbo-0125.
@@
 Due to the monetary costs of GPT-3.5-turbo-0125, we execute ComGPT twice for each starting seed node to handle GPT hallucinations.
-Evaluation metrics
-The evaluation metrics Fscore (Powers 2011) and Jaccard (Jaccard 1912) are used to evaluate the detected communities.
-For an overlapped seed node, all the communities it belongs to are integrated as a real local community.
-Fscore is the harmonic mean of precision and recall.
+The code is released on the author’s website [^1] https://github.com/cassius0301/ComGPT.
+III-A4 Evaluation metrics
+The evaluation metrics Jaccard [35] and Fscore [36] are used to evaluate the detected communities. Their formulas are as follows.
+$\text{Jaccard}=\frac{\left|C_{\text{Found}}\cap C_{\text{True}}\right|}{\left|C_{\text{Found}}\cup C_{\text{True}}\right|},$ (2)
+$\text{Fscore}=\frac{2\cdot\text{Precision}\cdot\text{Recall}}{\text{Precision}+\text{Recall}},$ (3)
+$\text{Recall}=\frac{\left|C_{\text{Found}}\cap C_{\text{True}}\right|}{\left|C_{\text{True}}\right|},$ (4)
+$\text{Precision}=\frac{\left|C_{\text{Found}}\cap C_{\text{True}}\right|}{\left|C_{\text{Found}}\right|},$ (5)
+where $C_{\text{Found}}$ is the local community that the algorithm discovers, and $C_{\text{True}}$ is the real local community to which the given starting node belongs.
 The values of Fscore and Jaccard are between 0 and 1.
 The larger value implies a better algorithm performance.
-Experimental results
-Results on networks
-For the Dolphins, Polbooks, and Football datasets, all node are selected as seed nodes.
-For the Amazon and DBLP datasets, we randomly select 200 nodes from the top 5,000 communities as seed nodes.
-The average Fscore and Jaccard of all seed nodes is calculated.
-The detailed experimental results are shown in Table 1.
-Table 1 shows that ComGPT outperforms the M method, R method, DMF_M, DMF_R, LCDNN, and LS on most datasets.
-Among all comparison methods, ComGPT exhibits the best performance on the Amazon and DBLP datasets and ranks second or third on the other three datasets.
-Although LS achieves the best performance on the Dolphins and Polbooks datasets, its performance on the remaining three datasets is the lowest among all comparison methods.
-Overall, ComGPT is superior to LS.
-ComGPT is superior to other methods because GPT-3.5-turbo has excellent comprehension to find the suitable node based on the NSG prompt.
-On the Football dataset, ComGPT performs worse than DMF_M, DMF_R, and LCDNN.
-It is because Football contains several tightly interconnected communities, which leads GPT-3.5-turbo to erroneously identify them as a single community.
+For an overlapped seed node, all the communities it belongs to are integrated as a real local community.
+III-B Experimental results
+TABLE III: Experimental results on five datasets.
+On each dataset, the highest Fscore (Jaccard) is highlighted in bold, the second highest one is highlighted by underline, and the third highest one is highlighted with a superscript asterisk.
+Dataset Metrics M R Louvain DMF_M DMF_R Leiden LCDNN LS FLMIG ComGPT
+Dolphins Jaccard 0.3370 0.3128 0.4705 0.6270 0.6320 0.5307 0.6438* 0.7735 0.5359 0.7035
+Fscore 0.4673 0.4478 0.5872 0.7254 0.7241 0.6320 0.7558* 0.8385 0.6517 0.7684
+Polbooks Jaccard 0.4266 0.3986 0.6403 0.6449 0.6437 0.6374 0.6673 0.6919 0.6336 0.6468*
+Fscore 0.5184 0.4960 0.7229 0.7294 0.7285 0.7237 0.7463 0.7764 0.7191 0.7320*
+Football Jaccard 0.6117 0.6134 0.8516* 0.8546 0.8579 0.7762 0.7547 0.2406 0.5596 0.7571
+Fscore 0.6706 0.6699 0.8890 0.8881* 0.8893 0.8361 0.8255 0.3593 0.6798 0.8103
+Amazon Jaccard 0.6961 0.6771 0.1917 0.7633 0.7686 0.4741 0.7665* 0.6506 0.0479 0.7734
+Fscore 0.7706 0.7554 0.2499 0.8353 0.8361* 0.5940 0.8372 0.7420 0.0642 0.8421
+DBLP Jaccard 0.3239 0.3207* 0.1786 0.2945 0.2928 0.2076 0.2208 0.0650 0.0074 0.3259
+Fscore 0.4185 0.4185 0.2657 0.4032 0.4009 0.3187 0.3152 0.0850 0.0144 0.4359
+III-B1 Results on networks
+For the Dolphins, Polbooks, and Football datasets, each node is taken as the starting node to obtain the local community.
+For the Amazon and DBLP datasets, we randomly select 200 nodes from the top 5,000 communities as the starting nodes.
+The average values of Fscore and Jaccard for the starting nodes are shown in Table III.
+Table III shows that ComGPT exhibits the best performance on the Amazon and DBLP datasets, ranking second and third places on the Dolphins and Polbooks datasets, respectively.
+On the Football dataset, ComGPT does not perform well.
+The reasons for the inconsistent performance of ComGPT across datasets are explained as follows.
+ComGPT achieves the highest performance on Amazon and DBLP.
+As shown in Table II, these datasets have low mixing parameters (0.1216 and 0.2147) and high modularity (0.44), indicating distinct community structures.
+This structure enables GPT-3.5-turbo to easily distinguish nodes inside and outside communities, thus achieving optimal results.
+For Polbooks and Dolphins, the modularity is lower than that of Amazon and DBLP, indicating weaker community structures.
+This increases the difficulty for GPT-3.5-turbo to accurately identify community members, resulting in some misidentified nodes and preventing ComGPT from achieving optimal performance.
+For Football, it has a high mixing parameter (0.3638), meaning numerous inter-community connections.
+These connections may cause GPT-3.5-turbo to merge multiple real communities into one, leading to ComGPT’s poor performance.
+LS performs well on datasets like Dolphins and Polbooks because it can accurately identify local leaders.
+This reliance on local leaders becomes a weakness in networks with strong inter-community connections, such as the football dataset.
+In these cases, a node identified as a local leader might actually be a follower of a more global leader, causing the algorithm to perform poorly.
+Louvain and its improved versions, such as Leiden and FLMIG, underperform on Amazon and DBLP datasets, as it tends to merge multiple distinct real communities into a single large community.
 Except for the DBLP dataset, DMF_M, DMF_R, and LCDNN perform better than the M method and R method.
-The reason why DMF_M, DMF_R, and LCDNN outperform the M method and R method on most datasets is that the latter two methods maximize local modularity ($M$ and $R$), making them sensitive to the choice of seed nodes, known as seed-dependent problem.
+The reason is that the latter two methods maximize local modularity ($M$ and $R$), making them sensitive to the choice of seed nodes, known as seed-dependent problem.
 In contrast, DMF_M, DMF_R, and LCDNN implement various strategies to alleviate this problem.
-Specifically, DMF_M, and DMF_R, utilize membership functions to select nodes with large degrees.
-LCDNN expands the community based on the fuzzy relations between nodes and NGC nodes.
-Ablation study
-Method GPT SK NSG NSU Fscore Jaccard
-1 NO NO NO NO 0.4666 0.3352
-2 YES NO YES YES 0.3711 0.2414
-3 YES YES NO YES 0.4855 0.3299
-4 YES YES YES NO 0.7134 0.6307
-ComGPT YES YES YES YES 0.7688 0.7009
-Table 2: Results of ablation experiment on the Dolphins dataset.
-To explore the necessity of GPT-3.5-turbo, SK (Supplementary knowledge in graph coding), NSG, and NSU prompts in ComGPT, we conduct ablation experiments on the Dolphins dataset with 50 randomly selected seed nodes.
+Specifically, DMF_M and DMF_R utilize membership functions to select nodes with large degrees, while LCDNN expands the community based on the fuzzy relations between nodes and their NGC nodes.
+ComGPT aims to explore the application of LLMs to local community detection.
+The baselines are mature methods, optimized and tuned for excellent performance.
+In contrast, ComGPT, an emerging exploratory method based on LLMs, is still in its early stages and requires further optimization to reach its best performance.
+Its value lies in revealing the potential of LLMs in local community detection.
+III-B2 Ablation study
+(a) Dolphins
+(b) Polbooks
+Figure 6: Ablation study.
+To explore the necessity of GPT-3.5-turbo, SK (Supplementary knowledge in graph coding), NSG, and NSU prompts in ComGPT, we conduct ablation experiments on the Dolphins and Polbooks datasets with 50 randomly selected seed nodes.
 Four simplified versions of ComGPT are constructed as follows: 1) Method 1 without GPT-3.5-turbo.
 The first node in $PN$ is selected for community expansion until $PN$ is empty; 2) Method 2 without SK; 3) Method 3 without NSG prompt; and 4) Method 4 without NSU prompt.
-When calculating Fscore and Jaccard, it is necessary to obtain the intersection between the detected community, which may include duplicate nodes, and the true community.
+The simplified version of ComGPT may add duplicate nodes to the detected communities.
+When calculating Fscore and Jaccard, it is necessary to obtain the intersection between the detected community and the true community.
 Here, duplicate nodes in the detected community are regarded as a single node.
+Additionally, retain the duplicate nodes in the detected community.
 To prevent cyclic selection of duplicate nodes, community expansion is terminated once the number of duplicate nodes exceeds one-third of the number of nodes in the community.
-Table 2 shows the experimental results of Methods 1, 2, 3, 4, and ComGPT.
-Overall, ComGPT outperforms Methods 1, 2, 3, and 4, indicating that GPT-3.5-turbo, SK, NSG, and NSU prompt play a crucial role in ComGPT, respectively.
-ComGPT is better than Method 1 in terms of Fscore and Jaccard, which demonstrates the effectiveness and importance of GPT-3.5-turbo.
+Figure 6 shows the results of Methods 1, 2, 3, 4, and ComGPT.
+Overall, ComGPT outperforms Methods 1, 2, 3, and 4, indicating that GPT-3.5-turbo, SK, NSG, and NSU prompts play a crucial role in ComGPT.
+ComGPT is better than Method 1 in terms of Fscore and Jaccard, which demonstrates the importance of GPT-3.5-turbo.
 The reason ComGPT outperforms Method 2 is that its graph coding incorporates community knowledge.
@@
 ComGPT also outperforms Method 3 and 4, suggesting that NSG and NSU prompts can enhance its performance in local community detection.
-Prompt comparisons
+III-B3 Prompt comparisons
+TABLE IV: Description of the prompts.
 Prompts Description
@@
 BaG Let’s construct a graph with the nodes and edges first.
-Table 3: Description of the prompts.
- Prompts Dolphins Polbooks Football
-Zero-shot Jaccard 0.3299 0.2191 0.5829
-Fscore 0.4855 0.3490 0.6774
-Few-shot Jaccard 0.3584 0.2703 0.7145
-Fscore 0.5128 0.4078 0.7764
-CoT Jaccard 0.3250 0.2078 0.6168
-Fscore 0.4785 0.3347 0.7196
-BaG Jaccard 0.2772 0.1813 0.6050
-Fscore 0.4235 0.2988 0.7131
-NSG Jaccard 0.7009 0.6624 0.8045
-Fscore 0.7688 0.7418 0.8555
-Table 4: Results of different prompts.
-To test the performance of the NSG prompt within ComGPT, we replaced the NSG prompt with four typical prompt approaches: Zero-shot (Brown et al.
-2020), Few-shot (Brown et al.
-2020), CoT (Wei et al.
-2022), and BaG (Wang et al.
-2023) prompts, as shown in Table 3.
-These alternate versions of the algorithm serve as baseline models for comparison.
-We conducted comparative experiments by randomly selecting 50 seed nodes in the Polbooks, Dolphins, and Football datasets.
-The detected communities may contain duplicate nodes, and the metrics values are calculated in the same way as in the ablation study.
-Table 4 shows that the performance of ComGPT with NSG is better than that of ComGPT with Zero-shot, Few-shot, CoT, and BaG prompts.
-Specifically, on Football dataset, the ComGPT with CoT and BaG outperform the ComGPT with Zero-shot.
-The ComGPT with CoT and BaG is worse than ComGPT with Zero-shot on Dolphins and Polbooks datasets.
+(a) Dolphins
+(b) Polbooks
+Figure 7: Results of different prompts.
+To test the performance of the NSG prompt within ComGPT, we replaced the NSG prompt with four typical prompt approaches: Zero-shot [37], Few-shot [37], CoT [38], and BaG [22] prompts, as shown in Table IV.
+We conducted comparative experiments by randomly selecting 50 seed nodes in the Polbooks and Dolphins datasets.
+Detailed experimental results are shown in Figure 7.
+From these results, we derive the following conclusions:
+Example samples are somewhat helpful to GPT-3.5-turbo.
+Based on the experimental results of Zero-shot and Few-shot, adding sample examples led to performance improvements in terms of Fscore and Jaccard on the Dolphins and Polbooks datasets.
+This indicates that providing examples for the GPT-3.5-turbo is indeed effective.
+Appropriate examples can improve performance.
 Advanced prompts may have limited performance improvement or even counterproductive effects on complex graph reasoning tasks.
-For instance, advanced prompts like CoT increase the number of reasoning steps for GPT-3.5-turbo, which potentially raises the likelihood of errors in the final reasoning conclusion due to mistakes in one of the intermediate steps.
-The reason ComGPT with Few-shot outperforms ComGPT with Zero-shot on most datasets is that we provided some examples of how the task should be executed.
-These examples guide the model by showing the correct process needed to solve the task.
-On Dolphins, Polbooks, and Football datasets, ComGPT with NSG outperforms ComGPT with other prompts.
+For instance, advanced prompts like CoT can sometimes introduce more points of failure, as each step must be correct to ensure a successful final decision.
+It will increase the risk of error accumulation, especially in complex graph-based tasks.
+Consequently, advanced prompts may not always improve performance.
+Compared to other prompts, NSG brings the highest improvement.
+Our designed prompt NSG achieve the best performance in terms of Fscore and Jaccard on Dolphins and Polbooks datasets.
 This is because the design of CoT and BaG is oriented towards universal graph reasoning problems so that no relevant guidance or knowledge is utilized in dealing with local community detection.
 NSG is oriented to the field of local community detection and utilizes the guidance in local community detection.
-Effect of the number of potential nodes on ComGPT
+III-B4 Effect of NSG_1 and NSG_2 on ComGPT
+We explore the effect of NSG_1 and NSG_2 on ComGPT on the polbooks dataset.
+In ComGPT, the NSG_1 prompt handles the seed-dependent problem and free rider effect, while the NSG_2 prompt tackles community diffusion (Section II-D).
+To verify whether ComGPT alleviates these shortcomings, we remove NSG_1 and NSG_2 from ComGPT to obtain methods named “No1” and “No2”, respectively.
+Result of No1, No2, ComGPT, and M Method on Polbooks is shown in Figure 8.
+Figure 8 shows that ComGPT, which combines the M method and GPT-3.5-turbo, outperforms the M method alone, indicating that ComGPT mitigates these shortcomings faced by the M method.
+ComGPT’s superiority over No1 indicates that the first prompt mitigates seed-dependent problem and free-rider effect, while its superiority over No2 shows that the second prompt mitigates community diffusion.
+III-B5 Effect of the number of potential nodes on ComGPT
+Figure 8: Effect of NSG_1 and NSG_2 on ComGPT.
 Figure 9: Effect of $k$ on ComGPT.
@@
 The values of $k$ are 5, 7, 10, 13, and 15.
-Figure 4 shows the performance of ComGPT under different $k$.
-As the $k$ increases, the performance of ComGPT remains generally stable, though with some fluctuations.
+Figure 9 shows the result of ComGPT under different $k$.
+As $k$ increases, the performance of ComGPT remains generally stable, though with some fluctuations.
 Fluctuations are relatively low across different values, except for $k$ = 10, which exhibits higher variability due to the randomness of GPT-3.5-turbo output.
-To confirm this, we conduct repeated experiments for $k$ = 10 and the results are nearly identical to those obtained with other values of $k$.
+To confirm this, we repeat experiments for $k$ = 10 and the results are nearly identical to those obtained with other values of $k$.
 The above analysis suggests that the performance of ComGPT is less affected by the parameter $k$.
-Conclusion
+III-B6 Analysis of GPT-3.5-turbo API Runtime in ComGPT
+TABLE V: Execution Times on node 1 of the Football dataset.
+GPT-3.5-turbo Response Time Other Execution Time
+t(s) Ratio(%) t(s) Ratio(%)
+1 133.4220 99.9441 0.0745 0.0559
+2 285.5551 99.9755 0.0699 0.0245
+3 150.5789 99.9595 0.0609 0.0405
+4 298.0469 99.9805 0.0579 0.0195
+5 128.2163 99.9540 0.0589 0.0460
+Average 199.1678 99.9677 0.0644 0.0323
+To analyze the runtime of calling the GPT-3.5-turbo API in ComGPT, we selected node 1 of the Football network as the starting node.
+We record the execution time of ComGPT at this starting node over five runs, including the GPT-3.5-turbo response time, the ratio of GPT-3.5-turbo response time in ComGPT, the local computation time, the proportion of local computation time in ComGPT, and the average values of these results.
+The experimental results are shown in Table V.
+The experimental results show that the response time of GPT-3.5-turbo accounts for more than 99.9% of the ComGPT’s execution time, indicating that the overall execution efficiency is mainly limited by GPT-3.5-turbo’s response speed rather than other computational processes.
+ComGPT needs to call GPT-3.5-turbo and wait for its response.
+This process is influenced by various factors, such as network conditions and server load, making it difficult to predict the exact response time.
+III-B7 Scalability
+To evaluate the scalability of ComGPT, we randomly sample subgraphs with different sizes from the Amazon dataset and test ComGPT on these extracted subgraphs.
+We extract subgraphs containing 20%, 40%, 60%, 80%, and 100% of the nodes from the Amazon dataset, respectively.
+We randomly select 100 nodes as the starting nodes from each subgraph.
+The average values of Precision, Recall, Fscore, Jaccard, runtime, and number of tokens are shown in Table VI.
+Table VI shows that, as the percentage of nodes increases, the Recall, F-score, and Jaccard of ComGPT improve.
+This is because, as the number of nodes and edges in the network increases, the community structure becomes more obvious, and the communities detected by ComGPT gradually become more complete.
+Furthermore, as the percentage of nodes increases, we observe that the Precision fluctuates while the Recall consistently increases.
+This growth in Recall indicates that the size of the detected communities also increases with the network size, covering more ground-truth members.
+The fluctuation in Precision suggests that this expansion is not always perfectly accurate.
+The total runtime and number of tokens of ComGPT increase as the network grows.
+As the percentage of nodes increases, the average degree of the nodes in the network rises.
+Since the time complexity is proportional to the average degree, the runtime increases as the percentage of nodes increases.
+As the percentage of nodes increases, the size of the detected communities also increases, leading to a higher number of required API calls.
+TABLE VI: Scalability of ComGPT on the Amazon dataset.
+20% 40% 60% 80% 100%
+Precision 0.9930 0.9657 0.9609 0.9520 0.9332
+Recall 0.1497 0.3121 0.4779 0.6973 0.8601
+Fscore 0.2452 0.4322 0.5954 0.7729 0.8646
+Jaccard 0.1490 0.3038 0.4594 0.6719 0.8033
+Time (s) 28.63 80.98 84.95 181.97 352.07
+Tokens 383.14 1344.21 4026.44 11431.69 12386.43
+IV Related work
+In this section, we briefly review the related work on community detection and the application of Large Language Models (LLMs) in graph analysis.
+IV-A Community detection
+Community detection aims to identify groups of closely interconnected nodes.
+According to whether they utilize global information of the network, existing community detection methods are classified into two types: global community detection and local community detection.
+IV-A1 Global community detection
+Global community detection depends on the network’s global information to discover all communities.
+In recent years, researchers have investigated various algorithms [39], such as the Label Propagation Algorithm (LPA) [40], Non-negative Matrix Factorization (NMF) [41], random walk based techniques [42], and so on.
+Raghavan et al.
+introduced the Label Propagation Algorithm (LPA) [43].
+In the LPA, each node iteratively updates its label to the label with the highest frequency.
+Communities are divided based on the distribution of labels.
+Wang et al.
+developed Non-negative Matrix Factorization (NMF)-based algorithms tailored for undirected networks, directed networks, and compound networks [44].
+The algorithm determines the membership of nodes to communities by decomposing the adjacency matrix.
+Meo et al.
+proposed that edge weights enhance community detection and developed the WERW-Kpath [45].
+It consists of two steps: (1) selecting a source node based on the degree of the node and the number of nodes in the network, and (2) choosing an unvisited edge from current node with probability proportional to its weight.
+The above studies require complete network data to identify all communities.
+In contrast, our work focuses on determining the community of a specific node using only local information, thereby fundamentally distinguishing it from global studies.
+IV-A2 Local community detection
+Local community detection focuses on discovering the community to which the seed node belongs [46, 47], without involving the entire network.
+Compared to global community detection, it is more efficient and avoids unnecessary computational overhead.
+A variety of methods, like $k$-core decomposition [48] and seed expansion algorithms [24], are developed for detecting local communities.
+Seed expansion algorithms expand communities by selecting nodes with high scoring function values [9].
+For example, the R method [24] and M method [4] choose nodes based on the local modularity R and M, respectively.
+The seed expansion algorithms often face the seed-dependent problem [10], community diffusion, and free rider effect [11, 12, 13].
+To address the seed-dependent problem, researchers developed different strategies such as identifying core node sets [14, 49], finding the nearest nodes with greater centrality [15], alternating strategy of strong fusion and weak fusion [16], and designing membership functions with a bias towards nodes with large degrees [29].
+Differing from the above studies, our work alleviates the aforementioned shortcomings by utilizing LLMs.
+IV-B LLMs in graph analysis
+IV-B1 Application of LLMs in graph reasoning
+In recent years, researchers have begun to explore the potential of LLMs for applications in the areas of graph reasoning and graph machine learning.
+Wang et al.
+designed the NLGraph dataset for evaluating the graph reasoning capabilities of LLMs and demonstrated that LLMs do have graph reasoning abilities [22].
+Zhang et al.
+designed the Graph-ToolFormer framework to handle graph reasoning tasks [50].
+It enabled LLMs to self-guide using ChatGPT-enhanced prompts.
+Chen et al.
+utilized LLMs for the node classification task and proposed two approaches: LLMs-as-Enhancers and LLMs-as-Predictors [51].
+The former leverages LLMs to enhance the node’s text attributes with their massive knowledge and then generates predictions through GNNs.
+The latter attempts to directly employ LLMs as standalone predictors.
+Chen et al.
+proposed a new label-free pipeline LLM-GNN to leverage LLMs for annotation, providing training signals on GNN for further prediction [52].
+Specifically, LLMs are leveraged to annotate a small portion of nodes and then GNNs are trained on LLMs’ annotations to make predictions for the remaining large portion of nodes.
+IV-B2 Graph encoding in LLMs
+In the field of LLMs in graph machine learning research, scholars have focused on how to improve the ability of LLMs in processing graph data.
+Guo et al.
+emphasized the design and improvement of graph encoding techniques, which critically affect LLMs’ interpretation of input data and its subsequent output [53].
+Therefore, Adjacency and Incident [23] are designed to encode graph structured data into text for LLMs.
+Adjacency uses integer node encoding and parenthesis edge encoding.
+Incident uses integer node encoding and incident edge encoding.
+Existing graph coding methods primarily represent graph topology, which is often insufficient for specific tasks, such as community detection.
+Unlike existing approaches, our work introduces a novel graph encoding method tailored for the community detection task.
+IV-B3 Prompts in LLMs
+PROMPT engineering is an approach to increase model inputs through prompts used for specific tasks so that LLMs are better adapted to new tasks [54].
+The prompt is widely used in computer vision [55], recommender systems [51], and other fields.
+This approach does not require modification of model parameters, but rather significantly enhances the model’s ability to process downstream tasks based on task-specific prompts.
+Various prompts are designed for specific tasks to make LLMs better adapt to these tasks [54], such as the general CoT (Chain-of-Thought) prompt [38] and the BaG (Build a Graph) prompt [22].
+CoT aims to enable LLMs to solve complex problems through step-by-step reasoning.
+Its core idea is to mimic human thought processes by reasoning incrementally rather than directly generating the final answer.
+BaG encourages LLMs to map the textual descriptions of graphs and structures to grounded conceptual spaces before tackling the specific problem through a one-sentence instruction.
+However, these prompts do not incorporate the domain knowledge required for local community detection and are unsuitable for detecting community structures within networks.
+This paper designs prompts suitable for local community detection tasks, which are different from the above prompts.
+V Conclusion
 We explored using GPT-3.5-turbo to assist in seed expansion algorithm and developed a GPT-guided methodology, termed ComGPT.
@@
 Furthermore, the NSG prompt is designed to enhance LLMs’ understanding of community domain knowledge.
-Experimental results confirm that the designed graph encoding method and prompts help GPT-3.5-turbo understand the domain knowledge and detect local community structure, respectively.
+Experimental results confirm that the designed graph encoding method and prompts help GPT-3.5-turbo understand the domain knowledge and detect local community structure.
+Although ComGPT utilizes a local expansion strategy to minimize token consumption, the size of local communities in large-scale networks may be so large that the input information exceeds the token limits of LLMs.
+In the future, we will explore scalable local community detection and information compression methods, aiming to ensure that when the community size is large, the information input to LLMs remains within their token limits.
+Besides, to explore the potential of the proposed idea, we plan to extend it to other tasks for 2D/3D multi-media data [56, 57].
+Acknowledgment
+This work was supported by the National Natural Science Foundation of China [No.62572002, No.62272001 and No.62206004], Natural Science Foundation of Anhui Province of China [No.2508085MF159], Hefei Key Technology R&D “Champion-Based Selection” Project [No.2023SGJ011], and Guangdong Provincial Key Laboratory of Novel Security Intelligence Technologies [No.2022B1212010005].